### v1

# Gram-like matrix preserving extensions of noncommutative polynomials to sum of Hermitian Squares

Abstract.
Given a nonnegative noncommutative polynomial $f$, equivalently a sum of Hermitian squares (SOHS), there exists a positive semidefinite Gram matrix that encrypts all essential information of $f$. There are no available methods for extending a noncommutative polynomial to a SOHS keeping the Gram matrices unperturbed. As a remedy, we introduce an equally significant notion of Gram-like matrices and provide linear algebraic techniques to get the desired extensions. We further use positive semidefinite completion problem to get SOHS and provide criteria in terms of chordal graphs and 2-regular projective algebraic sets.
Keywords : Noncommutative polynomial, sum of Hermitian squares, chordal graph, Castelnuovo-Mumford regularity, Betti table, free resolution, matrix completion.
2020 Mathematics Subject Classification : 14P05, 14N25, 13D02, 05C50, 11E25, 90C90.
1. Introduction
Hilbert’s 17th problem states whether every multivariate polynomial that takes only nonnegative values over the reals can be represented as sum of squares (SOS) of polynomials. Hilbert, in 1888, proved that it is true only for quadratic forms, binary forms and ternary quatrics in the commutative case. In the noncommutative setup, Helton proved that any polynomial is nonnegative if and only if it is a sum of Hermitian squares (SOHS). The SOS and SOHS techniques are heavily used for polynomial optimization and convex optimization, in particular, in semidefinite programming (SDP). Many of the SOHS techniques are taken from real algebraic geometry and inspired from functional analysis.
The SOHS techniques apply across many different disciplines including control theory, probability and statistics. Out of many directions, hierarchy techniques are still very active areas of research in quantum information theory (cf. [12] and [22]) and have been proven useful in discriminating the quantum states. A large portion of the theory for noncommutative polynomials is being developed by Helton, McCullough, Klep and their collaborators (cf. [10], [17], [18], [20] and [21]). The book “Optimization of Polynomials in Non-commuting Variables” (cf. [9]) by Burgdorf, Klep and Povh is a beautiful introduction to the subject. Pablo Parillo has used SOS techniques in different SDP to produce different control theory oriented applications (cf. [23]). Blekherman and his collaborators’ work on SOS techniques provide fruitful insight about the fundamental geometric properties of some projective algebraic sets (cf. [8], [6] and [7]). Motivated by the theory of noncommutative polynomials, noncommutative functions have been developed over noncommutative sets and many researchers are working on the subject (cf. [1] and [25]).
This paper is aimed at a large audience with minimal background in the noncommuatative polynomials. Most of the concepts are self-contained. One of the key aspects that the reader will take away from this paper is the technique to produce ample SOHS. Moreover, these techniques will be fruitful in SDP and its applications.
This article blends different areas, such as linear algebra, graph theory, algebraic geometry and commutative algebra, with a focus on SOHS techniques. Given a noncommutative polynomial parts of which are SOHS, it is always possible to find a SOHS extension $\tilde{f}$ by adding suitable terms with $f$. But, in general, a Gram matrix of such $\tilde{f}$ doesn’t preserve the Gram matrices of the SOHS parts of $f$. In this paper, we define a notion called Gram-like matrix, which is equally significant to, but more flexible than, a Gram matrix. Exploiting the obtained flexibility and Sylvester’s criteria, we provide techniques to preserve any Gram-like matrix of the SOHS part of some well behaved $f$ inside a Gram-like matrix of $\tilde{f}$ as a leading principal submatrix. Furthermore, we prove that a SOHS completion $\tilde{f}$ of a quasi SOHS polynomial $f$ is possible if the specification graph and the subspace arrangement of $f$ is chordal and $2$-regular respectively. In that case, a Gram-like matrix of $\tilde{f}$ contains Gram-like matrices of the SOHS parts of $f$ as its principal submatrices. In fact, we show that this SOHS completion is an alternative elucidation of the celebrated positive semidefinite completion problem.
We now go through the chronology of the paper in a bit more detail. In Section 2, we recall some prerequisites regarding noncommutative polynomials. We discuss the notion of symmetric polynomials, SOHS polynomials and some basic properties of positive definite and positive semidefinite matrices. Following that, we rigorously discuss the notion of a Gram matrix and monomial vector associated to a symmetric polynomial and SOHS. We observe that depending on the polynomial, the size of the associated Gram matrix can be reduced without affecting the defining properties. Moreover, one can be carefree about the order of the constituent monomials of the monomial vector. We point out these drawbacks of a Gram matrix through several examples (cf. $(2)$ of Remark 2.13 and Example 2.14). We sort these things out formally (cf. Theorem 3.5 and Theorem 3.6) in Section 3 and refer to the associated matrix as Gram-like matrix (cf. Definition 3.7 and Definition 3.8), which is analogous to, but more flexible than, Gram matrix (cf. Remark 3.9 for a comparison between a Gram matrix and a Gram-like matrix). Moreover, it keeps the soul of a Gram matrix intact. We then ask the following question :
Question 1.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial.
  1. (a)
    What is the procedure to get a SOHS extension $\tilde{f}$ of $f$?
  2. (b)
    If $f$ has some SOHS parts, can it be extended to a SOHS $\tilde{f}$ keeping the gram-like matrices of the the SOHS parts of $f$ intact?
Theorem 3.14 answers part $(a)$ of Question 1, whereas Theorem 3.16, Theorem 3.29 and Theorem 3.40 answers part $(b)$ of Question 1.
We first answer this question in the most trivial case, namely, adding a positive constant to a SOHS gives a SOHS (cf. Theorem 3.10). But we do so by exploiting the flexibility of a Gram-like matrix. Followed by that, we prove the first main theorem of this article (cf. Theorem 3.16).
Theorem 1.1.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial, without any constant term, which can be expressed as a SOHS plus $r$ many terms $a_{1}\zeta_{1},a_{2}\zeta_{2},\ldots,a_{r}\zeta_{r}$. Assume that ${\rm rc}(\eta,i)\neq\zeta_{j}$, $1\leq j\leq r$, for any $i\geq 0$ and for any monomial $\eta$ of the SOHS part of $f$. Then
  1. (1)
    It is always possible to obtain a SOHS $\widetilde{f}$ by adding at most $12r+$ 1 terms with $f$ such that there exists a positive semidefinite Gram-like matrix of $\widetilde{f}$ which contains a positive semidefinite Gram-like matrix of the SOHS part of $f$ as a leading principal submatrix.
  2. (2)
    In fact, infinitely many such $\widetilde{f}$ can be obtained.
We further observe that one can even control the degree of $\widetilde{f}$ in the aforementioned scenario (cf. Theorem 3.18).
In subsection 3.3, we deal with Question 1 via graph theory. In fact, we look at positive semidefinite completion problem in the context of completion of a quasi SOHS polynomial (cf. Definition 3.28 for details). We prove the following theorem which describes a pattern made out of the known coefficients of a quasi SOHS polynomial that ensures a SOHS completion of the same (cf. Theorem 3.29). The pattern is given in terms of a graph associated to a partial Gram-like matrix of the given quasi SOHS polynomial.
Theorem 1.2.
Let $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ be a quasi SOHS polynomial w.r.t a representation $\mathbf{R}_{f}$, as in (10), along with conditions as in (11). Then any such $f$ has a SOHS completion $\overline{f}^{\;G_{f},\;W_{k}}$ w.r.t the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$ if and only if the specification graph of $G_{f}$ is chordal. Here $W_{k}$ is the column vector $\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}\end{pmatrix}^{t}$, $\zeta_{i}$, $1\leq i\leq k$, being the monomials appearing in $\mathbf{R}_{f}$.
Here a positive semidefinite Gram-like matrix of any such SOHS completion keeps the positive semidefinite Gram-like matrices of SOHS parts of the given quasi SOHS polynomial as its principal submatrices. We also prove a similar statement for Gram matrix as well, (cf. $(1)$ of Theorem 3.29).
In subsection 3.4, we again answer when a quasi SOHS polynomial can be completed to a SOHS but now the pattern of known and unknown entries is guided by a projective algebraic set. Moreover, the defining equations of this algebraic set are homogeneous square-free quadratic (commutative) monomials. The nobleness of this elucidation is that a SOHS completion of some partial symmetric noncommutative polynomial of any even degree is controlled by some homogeneous (commutative) monomials of degree $2$. To be more specific, we prove that $2$-regularity of such a projective algebraic set guarantees the SOHS completion. Furthermore, it is deeply related to Theorem 1.2 as well. Precisely, we obtain the following (cf. Theorem 3.40 for details):
Theorem 1.3.
Let $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ be a quasi SOHS polynomial w.r.t a representation $\mathbf{R}_{f}$, as in (10), along with conditions as in (11). Then the following are equivalent:
  1. (1)
    Any such $f$ has a SOHS completion $\overline{f}^{\;G_{f},\;W_{k}}$ w.r.t the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$.
  2. (2)
    The subspace arrangement $V_{G_{f}}(\subseteq\mathbb{P}^{k-1})$ of $G_{f}$ satisfies $P_{V_{G_{f}}}=\Sigma_{V_{G_{f}}}$.
  3. (3)
    The subspace arrangement $V_{G_{f}}(\subseteq\mathbb{P}^{k-1})$ of $G_{f}$ is $2$-regular.
Here $W_{k}$ is the column vector $\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}\end{pmatrix}^{t}$, $\zeta_{i}$, $1\leq i\leq k$, being the monomials appearing in $\mathbf{R}_{f}$.
We provide some examples to display simultaneous applications of Theorem 3.29 and Theorem 3.40. In these examples we use Macculay2 software to calculate regularity via Betti tables and free resolutions.
2. Background material
We introduce our notations and recall necessary definitions and results in this section.
2.1. Noncommutative polynomial and SOHS
Let $X_{1},X_{2},\dots,X_{n}$ be noncommuting variables. Generating all possible words, also referred as monomials, in these letters with finite length by concatenating the letters and adding $1$, the empty word, we get a free monoid. Let $\langle\underline{X}\rangle=\langle X_{1},X_{2},\dots,X_{n}\rangle$ be the free monoid generated by $X_{1},X_{2},\dots,X_{n}$. We take all possible finite linear combinations of words from $\langle\underline{X}\rangle$ with coefficients from $\mathbb{R}$ and we call them real noncommutative polynomials. Denote the set of all real noncommutative polynomials as
$\mathbb{R}\langle\underline{X}\rangle=\mathbb{R}\langle X_{1},X_{2},\dots,X_{n}\rangle=\Big{\{}\sum_{i=1}^{N}a_{i}W_{i}\mid\;a_{i}\in\mathbb{R},N\in\mathbb{N},W_{i}\in\langle\underline{X}\rangle\Big{\}}.$
$\mathbb{R}\langle\underline{X}\rangle$ forms a free algebra. We equip $\mathbb{R}\langle\underline{X}\rangle$ with the * operation
$\mathbb{R}\langle\underline{X}\rangle$ $\longrightarrow\mathbb{R}\langle\underline{X}\rangle$
$f$ $\longmapsto f^{*}.$
Under this map, $\mathbb{R}\cup\{\underline{X}\}$ gets fixed and for given
$W$ $=X_{1}X_{2}\dots X_{n-1}X_{n},$
$W^{*}$ $=X_{n}X_{n-1}\dots X_{2}X_{1}.$
By the degree of a monomial $W$, denoted by ${\rm deg}(W)$, we mean the length of $W$. For example, ${\rm deg}(X_{1}X_{2}^{3}X_{3})={\rm deg}(X_{1}X_{2}X_{2}X_{2}X_{3})=5$. By the degree of a noncommutative polynomial $f$, denoted by ${\rm deg}(f)$, we mean the degree of any one of its highest degree monomials.
We reserve the notation ${\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ to denote the set of all symmetric polynomials of $\mathbb{R}\langle\underline{X}\rangle$, that is,
${\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle=\{f\in\mathbb{R}\langle\underline{X}\rangle\mid\;f=f^{*}\}.$
Let $\langle\underline{X}\rangle_{d}$ be the set of all words in $\langle\underline{X}\rangle$ of at most degree $d$. Let $\mathbb{R}\langle\underline{X}\rangle_{d}$ be the set of all polynomials with monomials coming from $\langle\underline{X}\rangle_{d}$. The cardinality of $\langle\underline{X}\rangle_{d}$ is $s(d,n)=\sum_{i=0}^{d}n^{i}$, where $n$ is the number of noncommuting variables in $\langle\underline{X}\rangle$. Let $W_{s(d,n)}$ be the $s(d,n)\times 1$ column vector of the words in $\langle\underline{X}\rangle_{d}$ taken in the lexicographic order. Given a polynomial ${f}\in\mathbb{R}\langle X\rangle_{d}$, $\mathbf{f}\Big{(}=({f_{w}})_{w\in\mathbb{R}\langle X\rangle_{d}}\Big{)}$ be its coefficient vector. Every polynomial ${f}$ of degree $d$ can be expressed as $f=\sum_{w\in\langle\underline{X}\rangle_{d}}f_{w}w={\mathbf{f}}^{T}W_{s(d,n)}=W_{s(d,n)}^{*}\mathbf{f}$.
Definition 2.1.
For any $p\in\mathbb{R}\langle\underline{X}\rangle$, the noncommutative polynomial $p^{*}p$ is called a Hermitian square. A noncommutative polynomial $f\in\mathbb{R}\langle\underline{X}\rangle$ is said to be a sum of Hermitian squares (or SOHS in short) if it can be written as a sum of some Hermitian squares, i.e. there exists noncommutative polynomials $g_{1},g_{2},\dots,g_{k}\in\mathbb{R}\langle X\rangle$ for some $k\in\mathbb{N}$ such that
$f=g_{1}^{*}g_{1}+g_{2}^{*}g_{2}+\dots+g_{k}^{*}g_{k}.$
We now recall a notion that formalizes any right end part of a given monomial (cf. [20, Section 3, p. 745] or [9, Definition 2.5, p. 38]).
Definition 2.2.
The right chip function ${\rm rc}:\langle\underline{X}\rangle\times\mathbb{N}\rightarrow\langle\underline{X}\rangle$ is defined by
${\rm rc}(W_{1}W_{2}\cdots W_{n},i)=\left\{\begin{array}[]{ll}W_{n-i+1}W_{n-i+2}\cdots W_{n}&\mbox{if $1\leq i<n$};\\ \\ W_{1}W_{2}\cdots W_{n}&\mbox{if $i\geq n$};\\ \\ 1&\mbox{if $i=0$}.\end{array}\right.$
Here ${\rm deg}(W_{i})=1$ for all $1\leq i\leq n$.
Example 2.3.
Let $W=X_{1}^{2}X_{2}^{2}X_{3}X_{2}^{2}X_{1}$ be a noncommutative word. Then, ${\rm rc}(W,4)=X_{3}X_{2}^{2}X_{1}$, ${\rm rc}(W,8)={\rm rc}(W,9)=X_{1}^{2}X_{2}^{2}X_{3}X_{2}^{2}X_{1}$ and ${\rm rc}(W,0)=1$.
2.2. Some properties of positive definite and semidefinite matrices
There are many different equivalent definitions for a symmetric matrix to be positive semidefinite. Here we present one with the matrix minors. A minor of a square matrix is the determinant of a submatrix obtained by deleting same number of rows and columns. A principal submatrix of a square matrix is the one where the indices of the deleted rows and the indices of the deleted columns are same. A principal minor is the determinant of a principal submatrix. A leading principal submatrix is a principal submatrix where index of the one deleted row (or column respectively) implies all the rows (or column respectively) with higher indices are deleted. In other words, the leading principal submatrices are square matrices sitting at the left corners of the actual matrix. Determinant of a leading principal matrix is called a leading principal minor.
We now recall the characterisations of positive definite and positive semidefinite matrices in terms of their principal minors.
Theorem 2.4 (Sylvester’s criterion for positive definite matrices).
A real symmetric matrix is positive definite if all its leading principal minors are positive.
Proof.
See [2, §3.3, 3.15, p. 27]. ∎
Analogous version doesn’t hold for a positive semidefinite matrix, i.e, nonnegativity of all leading principal minors doesn’t guarantee the positive semidefiniteness of a real symmetric matrix. For example, both the leading principal minors of the real symmetric matrix
$\begin{pmatrix}0&0\\ 0&-1\end{pmatrix}$
are 00. But the matrix clearly is not positive semidefinite as its eigenvalues are 00 and $-1$ (cf. [3, §1.2, p. 7]). Therefore, we need some extra conditions in case of positive semidefinite matrices.
Theorem 2.5 (Generalized Sylvester’s criterion for positive semidefinite matrices).
A real symmetric matrix is positive semidefinite if all its principal minors are nonnegative.
Proof.
See [5, §1.2, p. 7]. ∎
The following well-known results are easy consequences of spectral decomposition of a real symmetric matrix and will be used often in the proofs of the results in Section 3. The following results are direct consequences of the spectral theorem for real symmetric matrices.
Lemma 2.6.
Let $G$ be a positive semidefinite matrix with real entries of size $k$. Then the matrix obtained from $G$ by first interchanging its $i$-th and $j$-th row (column respectively) followed by interchanging its $i$-th and $j$-th column (row respectively) is again a positive semidefinite matrix, for any $1\leq i,j\leq k$.
Lemma 2.7.
Let $A$ be a block diagonal matrix with $A_{1},A_{2},\ldots,A_{n}$ being its diagonal blocks. If each $A_{i}$ is a positive definite (respectively positive semidefinite) matrix, then so is $A$.
We end this subsection by recalling one more well-known property of a positive semidefinite matrix. We present a rudimentary argument, which is available in the literature, for the reader not familiar with this result.
Lemma 2.8.
Let $A$ be a $k\times k$ positive semidefinite matrix with real entries. Let $(i,i)$-th entry of $A$ be zero, for some $1\leq i\leq k$. Then the $i$-th column and the $i$-th row of $A$ is zero.
Proof.
Follows from Theorem 2.5. Indeed, suppose the contrary. Then there will be some $1\leq j\leq k$ such that $(i,j)$-th entry, and hence $(j,i)$-th entry of $A$ is nonzero and both are equal. Then, the determinant of the $2\times 2$ principal submatrix of $A$ formed by $(i,i)$-th entry, $(i,j)$-th entry, $(j,i)$-th entry and $(j,j)$-th entry is negative, which contradicts Theorem 2.5. ∎
2.3. Gram Matrix of a symmetric polynomial and a SOHS
Though it is a well known fact that any symmetric noncommutative polynomial has a symmetric matrix associated to it, known as Gram matrix, we still include it in a bit detail as we need this later.
Theorem 2.9.
Any $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ of degree $2d$ in $n$ variables has a representation $\mathbf{R}_{f}$ given as follows:
$\mathbf{R}_{f}\;:\;f=\sum_{i=1}^{s(d,n)}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq s(d,n)}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$ (1)
where
$\begin{split}&a_{i},b_{ij}\in\mathbb{R}\;;\\ &\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle,\;{\rm deg}(\zeta_{i})\leq d\text{\;for\;all\;}1\leq i\leq k\;;\\ &s(d,n)=\sum_{k=0}^{d}n^{k}\;.\end{split}$ (2)
Hence, there exists a $s(d,n)\times s(d,n)$ symmetric matrix $G_{f}$ such that
$f=W_{s(d,n)}^{\ast}\;G_{f}\;W_{s(d,n)},$ (3)
where $W_{s(d,n)}$ is the column vector of all monomials of degree at most $d$ taken in lexicographic order.
Proof.
Firstly, it is easy to note that any degree $2k$ monomial can be written as a product of two degree $k$ monomials. Now, let $a\eta$ be a term of $f$. As $f$ is chosen from ${\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$, then either $\eta^{\ast}=\eta$ or $a\eta^{\ast}$ must be another term of $f$. Hence, the first part of the assertion follows.
Let $1=\zeta_{1}<\zeta_{2}<\cdots<\zeta_{s(d,n)}=X_{n}^{d}$ be all such monomials in ascending order, where we denote the lexicographic order in the set of all monomials by $<$. Then we have,
$W_{s(d,n)}=\left(\begin{array}[]{@{}c}\begin{matrix}\zeta_{1}&\zeta_{2}&\cdots&\zeta_{s(d,n)}\end{matrix}\end{array}\right)^{t}.$
Choosing
$G_{f}=\begin{pmatrix}a_{1}&b_{12}&b_{13}&\cdots&b_{1s(d,n)}\\ b_{12}&a_{2}&b_{23}&\cdots&b_{2s(d,n)}\\ b_{13}&b_{23}&a_{3}&\cdots&b_{3s(d,n)}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ b_{1s(d,n)}&b_{2s(d,n)}&b_{3s(d,n)}&\cdots&a_{s(d,n)}\end{pmatrix}\;,$ (4)
we have the second part of the assertion. ∎
Definition 2.10.
Let $f$ be any symmetric noncommutative polynomial. The matrix $G_{f}$, as in (4), is called a Gram matrix of $f$. It is also called the Gram matrix associated with the representation $\mathbf{R}_{f}$, as in (1), along with conditions as in (2).
Remark 2.11.
  1. (1)
    It can be noted that given a degree $2d$ symmetric noncommutative polynomial, it doesn’t necessarily have unique representation. For example, consider $f=5X_{1}^{6}+8X_{2}^{4}$. Then $f$ has representations $5(X_{1}^{3})^{\ast}X_{1}^{3}+4(X_{2})^{\ast}X_{2}^{3}+4(X_{2}^{3})^{\ast}X_{2}$ and $5(X_{1}^{3})^{\ast}X_{1}^{3}+4(X_{2}^{2})^{\ast}X_{2}^{2}+4(X_{2}^{2})^{\ast}X_{2}^{2}$.
  2. (2)
    However, any symmetric homogeneous polynomial of degree $2d$ has a unique representation as all terms of such a polynomial are of degree $2d$ and hence can be written uniquely in terms of product of two degree $d$ monomials.
  3. (3)
    The moment we fix a representation of a degree $2d$ symmetric noncommutative polynomial, associated Gram matrix and monomial vector get fixed. Therefore, any symmetric homogeneous polynomial of degree $2d$ has a unique Gram matrix, as mentioned in [9, Proposition 1.19, p. 7].
It is easy to note that any SOHS is definitely a symmetric noncommutative polynomial. We want to look at the converse. The following theorem gives a criteria for a symmetric polynomial to be a SOHS in terms of its Gram matrix.
Theorem 2.12.
A symmetric polynomial $f$ of degree $2d$ in noncommuting variables is a SOHS if and only if there exists a positive semidefinite matrix $G_{f}$ such that
$f=W_{s(d,n)}^{\ast}\;G_{f}\;W_{s(d,n)},$ (5)
where $W_{s(d,n)}$ is the column vector consisting of all words in $\langle\underline{X}\rangle$ of degree $\leq d$ taken in lexicographic order.
Proof.
See [11, §2, p. 105-110] or [9, Proposition 1.16, p. 6]. ∎
Remark 2.13.
  1. (1)
    At this point it is worthwhile to mention that to be able to write a SOHS polynomial $f$ as $f=W_{s(d,n)}^{\ast}\;G\;W_{s(d,n)}$ for some symmetric matrix $G$ doesn’t guarantee the nonnegativity of the matrix $G$. For example, we consider the polynomial (cf. [4, Example 3.4, p. 3])
    $f=1+2X_{1}+X_{1}^{2}+X_{1}X_{2}^{2}+2X_{2}^{2}+X_{2}^{2}X_{1}+X_{2}X_{1}^{2}X_{2}+X_{2}^{4}.$
    Then $f$ is an SOHS as it can be decomposed as
    $f=(1+X_{1}+X_{2}^{2})^{\ast}(1+X_{1}+X_{2}^{2})+(X_{1}X_{2})^{\ast}(X_{1}X_{2}).$
    Taking
    $G_{f}=\begin{pmatrix}1&1&0&0&0&0&0\\ 1&1&0&0&0&0&1\\ 0&0&0&0&0&0&0\\ 0&0&0&1&0&0&0\\ 0&0&0&0&2&0&0\\ 0&0&0&0&0&0&0\\ 0&1&0&0&0&0&1\end{pmatrix}\;,$
    we have $f=W_{7}^{\ast}\;G_{f}\;W_{7}$, where $W_{7}$ is given as
    $W_{7}=\begin{pmatrix}1&X_{1}&X_{1}^{2}&X_{1}X_{2}&X_{2}&X_{2}X_{1}&X_{2}^{2}\end{pmatrix}^{t}\;.$
    It can be noted that the symmetric matrix $G$ is not positive semidefinite as determinant of the principal sub matrix
    $\begin{pmatrix}1&1&0&0\\ 1&1&0&1\\ 0&0&1&0\\ 0&1&0&1\end{pmatrix}\;,$
    obtained from $G_{f}$ by deleting third, fifth and sixth row and column, is $-1$. One can look at [20, Example 2.2, p. 742] or [9, Example 1.18, p. 6-7] for another such example.
  2. (2)
    To express a symmetric polynomial of degree $2d$ in the form as in (3), it is not always necessary to take all the monomials of degree at most $d$ in $W_{s(d,n)}$. For example, consider the polynomial $f$ as in (1). Then for
    $W_{5}=\begin{pmatrix}1&X_{1}&X_{1}X_{2}&X_{2}&X_{2}^{2}\end{pmatrix}^{t}\;,$
    and
    $G=\begin{pmatrix}1&1&0&0&0\\ 1&1&0&0&1\\ 0&0&1&0&0\\ 0&0&0&2&0\\ 0&1&0&0&1\end{pmatrix}\;,$
    we have $f=W_{5}^{\ast}\;G\;W_{5}$.
As the choice of the monomial vector $W_{s(d,n)}$ is very strict in Theorem 2.9 and Theorem 2.12, there are several techniques to throw out the unnecessary monomials and making the matrix of smaller size. For example Newton chip method, augmented Newton chip method etc (cf. [9, Section 2.3 & 2.4]) has been used to reduce the size of the column vector $W_{s(d,n)}$, as in (5), and hence the size of matrix $G_{f}$. Moreover, while putting the monomials in $W_{s(d,n)}$, usually lexicographic order is followed. But as long as enough monomials is being collected in the monomial vector to express a given SOHS $f$, the order doesn’t matter. We demonstrate this through the following example.
Example 2.14.
Take $f=X_{1}^{2}+4X_{1}X_{2}+4X_{2}X_{1}+16X_{2}^{2}$. Then, for $(W^{\prime}_{2})^{\ast}=\begin{pmatrix}X_{1}&X_{2}\end{pmatrix}$, we have $G^{\prime}=\begin{pmatrix}1&4\\ 4&16\end{pmatrix}$ such that $f=(W^{\prime}_{2})^{\ast}\;G^{\prime}\;W_{2}^{\prime}$. Now let’s change the order of the monomials. Take, $(W^{\prime\prime}_{2})^{\ast}=\begin{pmatrix}X_{2}&X_{1}\end{pmatrix}$, we have $G^{\prime\prime}=\begin{pmatrix}16&4\\ 4&1\end{pmatrix}$ such that $f=(W^{\prime\prime}_{2})^{\ast}\;G^{\prime\prime}\;W_{2}^{\prime\prime}$.
It can be noted that the matrix $G^{\prime\prime}$ is obtained from the positive semidefinite matrix $G^{\prime}$ by interchanging its rows and columns and still is remaining a positive semidefinite matrix. This is what is allowing us not to bother about the order of the monomials in $W_{2}$. Lemma 2.6 says that this is not at all surprising.
It is evident from part $(2)$ of Remark 2.13 and Example 2.14 that to be able to reduce the size of the monomial vector (and hence that of a Gram matrix) whenever possible and write the monomial entries in that monomial vector in any arbitrary order, while expressing a symmetric or SOHS polynomial in the required form, puts us in a resilient situation. Therefore, we formalize all of this in the upcoming section after putting things in a particular perspective.
3. Extending a noncommutative polynomial to a SOHS
In this section, we address the following question, which captures the main theme of this article.
Question 1.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial.
  1. (a)
    What is the procedure to get a SOHS extension $\tilde{f}$ of $f$?
  2. (b)
    If $f$ has some SOHS parts, can it be extended to a SOHS $\tilde{f}$ keeping the gram-like matrices of the the SOHS parts of $f$ intact?
We now prepare the ground for defining Gram-like matrices, which will help us answering Question 1.
3.1. Gram-like matrices for symmetric and SOHS noncommutative polynomials
Recall that Theorem 2.12 provides a way of getting a SOHS polynomial of degree $2d$ in $n$ variables whenever we have a $s(d,n)\times s(d,n)$ positive semidefinite matrix. We observe that a positive semidefinite matrix of arbitrary size also produces a SOHS of correct degree.
Proposition 3.1.
Let $G$ be a $k\times k$ positive semidefinite matrix and $W_{k}$ be a column vector of size $k$ with arbitrary monomial entries (not necessarily in lexicographical order). Then
  1. (1)
    The polynomial $W_{k}^{\ast}\;G\;W_{k}$ is a SOHS polynomial.
  2. (2)
    Moreover, if no row (or column respectively) of $G$ is $\begin{pmatrix}0&0&\cdots&0\end{pmatrix}$ (or $\begin{pmatrix}0&0&\cdots&0\end{pmatrix}^{t}$ respectively), then degree of the polynomial $W_{k}^{\ast}\;G\;W_{k}$ is twice the degree of any of the highest degree monomials in $W_{k}$.
Proof.
Let $\zeta_{i}$, $1\leq i\leq k$, be any $k$ many monomials taken from $\mathbb{R}\langle\underline{X}\rangle$. Let
$W_{k}=\begin{pmatrix}\zeta_{1}&\zeta_{2}&\cdots&\zeta_{k}\end{pmatrix}^{t}\;.$
There exist some $i_{0}$ between $1$ and $k$ such that for all $1\leq i\leq k$,
${\rm deg}(\zeta_{i})\leq{\rm deg}(\zeta_{i_{0}})=d\;(\text{say}).$
let us now form $W_{s(d,n)}^{\text{unordered}}$ by adding remaining $s(d,n)-k$ many monomials $\eta_{j}$s, $1\leq j\leq s(d,n)-k$ from $\mathbb{R}\langle\underline{X}\rangle$ of degree at most $d$ with $W_{k}$, i.e,
$W_{s(d,n)}^{\text{unordered}}=\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}&\eta_{1}&\cdots&\eta_{s(d,n)-k}\end{pmatrix}^{t}\;.$
Let $G^{\prime}$ be the $s(d,n)\times s(d,n)$ matrix given as follows:
$G^{\prime}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\mbox{\Large 0}\end{array}\right)\;.$
We then have,
$W_{k}^{\ast}\;G\;W_{k}={W_{s(d,n)}^{\text{unordered}}}^{\ast}\;G^{\prime}\;W_{s(d,n)}^{\text{unordered}}.$ (6)
Clearly, there exists a permutation $\sigma$ in the symmetric group $S_{s(d,n)}$ on $s(d,n)$ many symbols such that entries of the new column vector $W_{s(d,n)}$ (say), obtained by applying $\sigma$ to the rows of $W_{s(d,n)}^{\text{unordered}}$, is in lexicographic order. By $G^{\prime\prime}$ we denote the matrix obtained from $G^{\prime}$ by permuting its rows and columns by $\sigma$. Then we have,
${W_{s(d,n)}^{\text{unordered}}}^{\ast}\;G^{\prime}\;W_{s(d,n)}^{\text{unordered}}=W_{s(d,n)}^{\ast}\;G^{\prime\prime}\;W_{s(d,n)}.$ (7)
Clearly, the matrix $G^{\prime}$ is a positive semidefinite matrix as $G$ is so and by Lemma 2.7. Therefore, $G^{\prime\prime}$ is also a positive semidefinite matrix by Lemma 2.6 and the fact that $\sigma$ is composition of transpositions. Hence the first part of the assertion follows from Theorem 2.12, (6) and (7).
Let $\zeta_{i_{0}}$ be one of the monomials in $W_{k}$ of maximum degree. Then, by hypothesis, the $i$-th column and $i$-th row is nonzero. Therefore, by Lemma 2.8, the $(i_{0},i_{0})$-th entry of $G$ is positive, say $a_{i_{0}}$. Then, $f$ contains the term $a_{i_{0}}\zeta^{\ast}\zeta$ of degree $2{\rm deg}(\zeta_{i_{0}})$. So, ${\rm deg}(f)\geq 2{\rm deg}(\zeta_{i_{0}})$. The inequality ${\rm deg}(f)\leq 2{\rm deg}(\zeta_{i_{0}})$ is obvious as $W_{d}$ contains no monomial of degree bigger than ${\rm deg}(\zeta_{i_{0}})$. Hence, the second part of the assertion follows. ∎
Corollary 3.2.
Let $G$ be any $k\times k$ symmetric matrix and $W_{k}$ be any column vector of size $k$ with arbitrary monomial entries (not necessarily in lexicographical order). Then the polynomial $W_{k}^{\ast}\;G\;W_{k}$ is a symmetric polynomial.
Proof.
Proof goes along the same line of the proof of the first part of the assertion of Proposition 3.1. ∎
Remark 3.3.
  1. (1)
    For a positive definite matrix $G$, every diagonal entry, being a principal minor of $G$, is positive. Therefore, no column or row of $G$ is zero by Lemma 2.8. So, the degree of the SOHS $W_{k}^{\ast}\;G\;W_{k}$ is twice the degree of any of the highest degree monomials in $W_{k}$.
  2. (2)
    Assertion $(2)$ of Proposition 3.1 doesn’t hold for a symmetric matrix in general. Let us take
    $W_{4}:=\begin{pmatrix}1&X_{1}&X_{2}&X_{2}^{2}X_{3}^{2}\end{pmatrix}^{t}\;,$
    and
    $G:=\begin{pmatrix}5&0&0&0\\ 0&1&0&3\\ 0&0&1&0\\ 0&3&0&0\end{pmatrix}\;.$
    Then the polynomial $f=W_{4}^{\ast}\;G\;W_{4}=X_{1}^{2}+X_{2}^{2}+3X_{1}X_{2}^{2}X_{3}^{2}+3X_{3}^{2}X_{2}^{2}X_{1}+5$ is a symmetric polynomial of degree $5$ and the highest degree monomial $X_{2}^{2}X_{3}^{2}$ has degree $4$.
  3. (3)
    Recall that Remark 2.13 provides an example of the fact that writing a SOHS polynomial in the form $W_{k}^{\ast}\;G\;W_{k}$ doesn’t guarantee the nonnegativity of $G$. Here we provide a simpler example incorporating the idea of the example provided in $(2)$ of Remark 3.3. Clearly, $f=2X_{1}X_{2}^{2}X_{1}$ is a SOHS as $f=(\sqrt{2}X_{2}X_{1})^{\ast}(\sqrt{2}X_{2}X_{1})$. For $W_{2}=\begin{pmatrix}1&X_{1}X_{2}^{2}X_{1}\end{pmatrix}^{t}$ and $G=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, $f=W_{2}^{\ast}\;G\;W_{2}$. But clearly $G$ is not positive semidefinite.
Let $W_{k}$ be a monomial vector with degree of its monomial entries being at most $d$ and $G$ be a $k\times k$ symmetric matrix. Then, as observed in $(2)$ of Remark 3.3, the symmetric polynomial $W_{k}^{\ast}\;G\;W_{k}$ might have degree much lesser than $2d$, unlike the positive semidefinite case. But it is not so under some mild conditions on the symmetric matrix involved.
Proposition 3.4.
Let $G$ be a symmetric matrix with positive diagonal entries and $W_{k}$ be a column vector of monomials from $\mathbb{R}\langle\underline{X}\rangle$. Then the polynomial $f:=W_{k}^{\ast}\;G\;W_{k}$ is a symmetric polynomial of degree $2d$, $d$ being the degree of any of the highest degree monomial entries of $W_{k}$.
Proof.
Firstly, such an $f$ is a symmetric polynomial by Corollary 3.2. Now let $\zeta_{i}$ be a monomial entry of $W_{k}$ having highest degree, say $d$, among all the entries of $W_{k}$ for some $1\leq i\leq k$. Let $a_{i}(>0)$ be the $(i,i)$-th entry of $G$. Then $a_{i}\zeta^{\ast}\zeta$ is a term of $f$ and hence ${\rm deg}(f)\geq 2d$. The other inequality is obvious as no monomial entry of $W_{k}$ has degree more than $d$. ∎
We now record the upshot of the findings of this subsection through a couple of theorems.
Theorem 3.5.
  1. (1)
    Let $G$ be a $k\times k$ positive semidefinite matrix with no row (or column) being zero and $W_{k}$ be a column vector of size $k$ with arbitrary monomial entries of degree at most $d$. Then the polynomial $f=W_{k}^{\ast}\;G\;W_{k}$ is a SOHS polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$ given as follows :
    $\mathbf{R}_{f}\;:\;f=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$ (8)
    where
    $a_{i},b_{ij}\in\mathbb{R},\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle,\;{\rm deg}(\zeta_{i})\leq d\text{\;for\;all\;}1\leq i<j\leq k\;,$ (9a)
[…]
    for any $\begin{split}&\text{for\;any\;}K=\{i_{0},i_{1},\ldots,i_{\mid K\mid}\}\subset\{1,2,\ldots,k\}\text{\;with\;}i_{0}<i_{1}<\cdots<i_{\mid K\mid}\\ &\text{the\;matrix\;}\begin{pmatrix}a_{i_{1}}&b_{i_{1}i_{2}}&\cdots&b_{i_{1}i_{\mid K\mid}}\\ b_{i_{1}i_{2}}&a_{i_{2}}&\cdots&b_{i_{2}i_{\mid K\mid}}\\ \vdots&\vdots&\ddots&\vdots\\ b_{i_{1}i_{\mid K\mid}}&b_{i_{2}i_{\mid K\mid}}&\cdots&a_{i_{\mid K\mid}}\end{pmatrix}\text{is\; positive\;semidefinite\;.}\end{split}$ (9c)
  2. (2)
    Conversely, let $f$ be a SOHS polynomial of degree $2d$ with a representation $\mathbb{R}_{f}$, as in (8), satisfying the conditions as given in (9). Then there exists a $k\times k$ positive semidefinite matrix $G$ with no row (or column) being zero and a column vector $W_{k}$ of size $k$ with monomial entries of degree at most $d$ such that $f=W_{k}^{\ast}\;G\;W_{k}$.
Proof.
The condition as in (9b) ensures that no column or row are full zero, that is,
[…]
by Proposition 3.4 and Lemma 2.8. Moreover, the condition as in (9c) ensures that the matrix whose $(i,j)$-th entries are $b_{ij}$ for all $i\neq j$, and $(i,i)$-th entries are $a_{i}$ for all $i$, is positive semidefinite by Theorem 2.5. Now it follows from Proposition 3.1. ∎
Theorem 3.6.
  1. (1)
    Let $G$ be a $k\times k$ symmetric matrix with positive diagonal entries and $W_{k}$ be a column vector of size $k$ with arbitrary monomial entries of degree at most $d$. Then the polynomial $f=W_{k}^{\ast}\;G\;W_{k}$ is a symmetric polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$ given as follows :
    $\mathbf{R}_{f}\;:\;f=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$ (10)
    where
    $a_{i},b_{ij}\in\mathbb{R},\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle,\;{\rm deg}(\zeta_{i})\leq d\text{\;for\;all\;}1\leq i<j\leq k\;,$ (11a)
    $a_{i}>0\text{\;for\;all\;}1\leq i\leq k\;,$ (11b)
  2. (2)
    Conversely, let $f$ be a symmetric polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$, as in (10), satisfying the conditions as given in (11). Then there exists a $k\times k$ symmetric matrix $G$ with positive diagonal entries and a column vector $W_{k}$ of size $k$ with monomial entries of degree at most $d$ such that $f=W_{k}^{\ast}\;G\;W_{k}$.
Proof.
The condition as in (11b) ensures that no column or row are full zero. The assertion now follows from Proposition 3.4. ∎
Definition 3.7.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a symmetric polynomial of degree $2d$ in $n$ variables and $k$ be any positive integer satisfying $k\leq s(d,n)$. If there exist a $k\times k$ symmetric matrix $G$ with positive diagonal entries such that $f$ can be written in the form $W_{k}^{\ast}\;G\;W_{k}$ for some column vector $W_{k}$ of monomials of degree at most $d$, then $G$ is called a Gram-like matrix of $f$ (associated with $\mathbf{R}_{f}$, as in (10), along with conditions as in (11)).
Definition 3.8.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a SOHS polynomial of degree $2d$ in $n$ variables and $k$ be any positive integer satisfying $k\leq s(d,n)$. If there exists a $k\times k$ positive semidefinite matrix $G$ such that $f$ can be written in the form $W_{k}^{\ast}\;G\;W_{k}$ for some column vector $W_{k}$ of monomials of degree at most $d$, then $G$ is called a positive semidefinite Gram-like matrix of $f$ (associated with $\mathbf{R}_{f}$, as in (8), along with conditions as in (9)).
Remark 3.9.
  1. (1)
    The difference between a Gram matrix and a Gram-like matrix of a polynomial can be seen through the following example. A positive semidefinite Gram-like matrix of the polynomial $f=X_{1}^{2}+4X_{1}X_{2}+4X_{2}X_{1}+16X_{2}^{2}$, as in Example 2.14, is $G^{\prime\prime}=\begin{pmatrix}16&4\\ 4&1\end{pmatrix}$, whereas each of its positive semidefinite Gram matrix is of size $7$. One of its positive semidefinite Gram matrix $G_{f}$ is given as follows
    $G_{f}=\begin{pmatrix}0&0&0&0&0&0&0\\ 0&1&0&0&4&0&0\\ 0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0\\ 0&4&0&0&16&0&0\\ 0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0\end{pmatrix}\;.$
  2. (2)
    It can be noted from the proof of Proposition 3.1 that a positive semidefinite Gram-like matrix associated to a SOHS polynomial is equally important to its positive semidefinite Gram matrix as it encodes all the information of its Gram matrix. Moreover, it is easier to handle a Gram-like matrix as it usually has much lesser size than corresponding Gram matrix.
  3. (3)
    The motivation for working with a Gram-like matrix of a noncommutative polynomial $f$ instead of its Gram matrix is to avoid unnecessary monomials in the corresponding monomial vector, which is same as ignoring any full zero row (or column) from a Gram matrix. The transition from a Gram matrix to a Gram-like matrix can be interpreted through the lens of graph theory as well. To be specific, the sparsity pattern (cf. [24, §8.1, p. 329-330] for more details) of a Gram-like matrix is always connected, whereas it is not so for a Gram matrix in general.
3.2. Block Matrix Technique
In this subsection, we answer Question 1 in various setup using “block matrix technique” by exploiting the flexibility obtained through Theorem 3.5 and Theorem 3.6. In due course, we prevail this technique.
Let us begin with the most basic setup. Given any noncommutative SOHS $f$, it is easy to observe that $f+c$ is also a SOHS for any $c>0$. As a warm up, we again prove this incorporating the mentioned block matrix technique.
Theorem 3.10.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative SOHS polynomial. Then $f+c$ is again a SOHS polynomial for any $c>0$.
  1. (1)
    If $f$ has a constant term, there there exists a Gram-like matrix of $f+c$ that can be obtained from a Gram-like matrix of $f$ by changing only one diagonal entry.
  2. (2)
    If $f$ does not have a constant term, then there exists a Gram-like matrix of $f+c$ that contains a Gram-like matrix of $f$ as a leading principal submatrix.
Proof.
If $f$ is a SOHS, then so is $f+c$ for any $c>0$, as $f+c=f+(\sqrt{c})^{\ast}(\sqrt{c})$.
  1. (1)
    As $f$ is a SOHS of degree $2d$ (say), and it has a constant term, then by Theorem 3.5 there exists a positive semidefinite Gram-like matrix $G_{f}$ and a column vector $W_{k}$ of monomials of degree at most $d$ with $1$ as the $(i_{0},1)$-th entry such that $f=W_{k}^{\ast}\;G_{f}\;W_{k}$. We now consider the matrix $\widetilde{G_{f}}$, given as follows :
    $(\widetilde{G_{f}})_{ij}=\left\{\begin{array}[]{ll}(G_{f}){{}_{ij}}&\mbox{if $(i,j)\neq(i_{0},i_{0})$};\\ \\ (G_{f}){{}_{i_{0}i_{0}}}+c&\mbox{\text{otherwise}}.\end{array}\right.$ (12)
    Then clearly, as $c>0$, all the principal minors of $\widetilde{G_{f}}$ is greater than equal to that of $G_{f}$ and $G_{f}$ is a positive semidefinite matrix. Therefore, $\widetilde{G_{f}}$ is also so by Theorem 2.5. Moreover,
    $f+c=W_{k}^{\ast}\;\widetilde{G_{f}}\;W_{k},$ (13)
    where $\widetilde{G_{f}}$ is as in (12). So, the assertion follows from (13) and Theorem 3.5.
  2. (2)
    As $f$ is a SOHS of degree $2d$ (say), and it is not having any constant term, by Theorem 3.5 there exists a positive semidefinite Gram-like matrix $G_{f}$ and a column vector $W_{k}$ of monomials of degree at most $d$ with none of its entries being $1$ such that $f=W_{k}^{\ast}\;G_{f}\;W_{k}$. We now consider the matrix $\widetilde{G_{f}}$ of size $k+1$, given as follows :
    $\widetilde{G_{f}}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{f}\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}c\end{matrix}\end{array}\right)\;.$ (14)
    We also define $\widetilde{W_{k+2}}$ as follows.
    $\widetilde{W_{k+1}}:=\left(\begin{array}[]{@{}c}\begin{matrix}W_{k}\end{matrix}\\ \begin{matrix}1\end{matrix}\end{array}\right)\;.$ (15)
    Then, clearly we have
    $f+c\;=\;\widetilde{W_{k+1}}^{\ast}\;\widetilde{G_{f}}\;\widetilde{W_{k+1}},$ (16)
    where $\widetilde{W_{k+1}}$ and $\widetilde{G_{f}}$ are as in (15) and (14) respectively. Now as $c>0$ and $G_{f}$ is a positive semidefinite matrix, so is $\widetilde{G_{f}}$ by Lemma 2.7. Hence, the assertion follows from (16) and Theorem 3.5.
Corollary 3.11.
Let $A$ be a positive semidefinite matrix of size $n$. Then for any positive real numbers $b_{1},b_{2},\cdots b_{n}$, the matrix $\widetilde{A}$ given by
$\widetilde{A}_{ij}=\left\{\begin{array}[]{ll}A_{ij}&\mbox{if $i\neq j$};\\ \\ A_{ii}+b_{i}&\mbox{\text{otherwise}}.\end{array}\right.$
Proof.
Follows from the repeated applications of proof of the part $(1)$ of Theorem 3.10 and Lemma 2.6. ∎
Example 3.12.
Consider $f=X_{1}^{2}+X_{2}^{2}+3X_{1}X_{2}\in\mathbb{R}\langle X_{1},X_{2}\rangle$. Clearly, $f$ is not a SOHS as it doesn’t even lie in ${\rm Sym}\;\mathbb{R}\langle X_{1},X_{2}\rangle$. Now we consider the polynomials $\widetilde{f}_{i}$ s, by adding some terms to $f$, given as follows :
$\begin{split}\widetilde{f}_{1}&=f+3X_{2}X_{1}+4X_{2}X_{1}^{2}X_{2}+1=X_{1}^{2}+X_{2}^{2}+3X_{1}X_{2}+3X_{2}X_{1}+4X_{2}X_{1}^{2}X_{2}+1,\\ \widetilde{f}_{2}&=f+3X_{2}X_{1}+X_{2}X_{1}^{2}X_{2}+4=X_{1}^{2}+X_{2}^{2}+3X_{1}X_{2}+3X_{2}X_{1}+X_{2}X_{1}^{2}X_{2}+4,\\ \widetilde{f}_{3}&=f+3X_{2}X_{1}+9X_{2}X_{1}^{2}X_{2}+1=X_{1}^{2}+X_{2}^{2}+3X_{1}X_{2}+3X_{2}X_{1}+9X_{2}X_{1}^{2}X_{2}+1,\\ \widetilde{f}_{4}&=f+3X_{2}X_{1}+X_{2}X_{1}^{2}X_{2}+9=X_{1}^{2}+X_{2}^{2}+3X_{1}X_{2}+3X_{2}X_{1}+X_{2}X_{1}^{2}X_{2}+9.\end{split}$
The SOHS extensions $\widetilde{f}_{1}$ and $\widetilde{f}_{2}$ of $f$ can be decomposed as follows:
$\begin{split}\widetilde{f}_{1}&=(X_{1}+X_{2})^{\ast}(X_{1}+X_{2})+(2X_{1}X_{2}+1)^{\ast}(2X_{1}X_{2}+1),\\ \widetilde{f}_{2}&=(X_{1}+X_{2})^{\ast}(X_{1}+X_{2})+(X_{1}X_{2}+2)^{\ast}(X_{1}X_{2}+2).\end{split}$ (17)
Similarly, the SOHS extensions $\widetilde{f}_{3}$ and $\widetilde{f}_{4}$ of $f$ has following decompositions:
$\begin{split}\widetilde{f}_{3}&=X_{1}^{\ast}X_{1}+X_{2}^{\ast}X_{2}+(3X_{1}X_{2}+1)^{\ast}(3X_{1}X_{2}+1),\\ \widetilde{f}_{4}&=X_{1}^{\ast}X_{1}+X_{2}^{\ast}X_{2}+(X_{1}X_{2}+3)^{\ast}(X_{1}X_{2}+3).\end{split}$ (18)
Remark 3.13.
  1. (1)
    It can be noted that in the first set of decompositions, as given in (17), pre existing SOHS part of $f$ is getting affected. In contrast, in the decompositions, as given in (18), the same is remaining intact.
  2. (2)
    The decompositions, as in (18), are more natural in the sense that these do not depend on any sort of relation between the SOHS part and non-SOHS part of $f$, and therefore have the potential to provide positive semidefinite Gram-like matrices of $\widetilde{f}_{i}$ s, $i=3,4$, incorporating positive semidefinite Gram-like matrices of $f$.
We record these observations in a proper setup through the following theorems.
Theorem 3.14.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial which can be expressed as a SOHS plus $n$ many terms. Then it is always possible to obtain a SOHS $\widetilde{f}$ by adding at most $12n+$ 1 many terms with $f$.
Proof.
Let ${\rm deg}(f)=d$. By hypothesis, $f$ can be written as follows:
$f=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+\sum_{j=1}^{n}a_{j}\zeta_{j},$
where $a\in\mathbb{R}$, $g_{i}$s and $\zeta$ lie in $\mathbb{R}\langle\underline{X}\rangle$, ${\rm deg}(g_{i})\leq\tfrac{d}{2}$ if $d$ is even or ${\rm deg}(g_{i})\leq\tfrac{d-1}{2}$ if $d$ is odd, for all $1\leq i\leq m$, and $1\leq{\rm deg}(\zeta_{j})\leq d$ for all $1\leq j\leq n$. We define a polynomial $h$, consisting of $12n+$ 1 terms, as follows:
$h:=\sum_{j=1}^{n}a_{j}\zeta_{j}^{\ast}+\sum_{j=1}^{n}\zeta_{j}^{\ast}\zeta_{j}+(\sum_{j=1}^{n}a_{j}^{2}).$
Then $\widetilde{f}(:=f+h)$ is a SOHS as it has the following decomposition:
$\widetilde{f}=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+\sum_{j=1}^{n}(\zeta_{j}+a_{j})^{\ast}(\zeta_{j}+a_{j}).$
We now aim to provide a positive semidefinite Gram-like matrix for the extended SOHS $\widetilde{f}$ in terms of a positive semidefinite Gram-like matrix of the partial SOHS $f$. For that we put some conditions on the non-SOHS part of $f$ and then use block-matrix technique, as introduced in the proof of Part - 2 of Theorem 3.10. We first prove it assuming that nonconstant part of $f$ is made out of a single term.
Theorem 3.15.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial, without any constant term, which can be expressed as a SOHS plus another term $a\zeta$. Assume that ${\rm rc}(\eta,i)\neq\zeta$ for any $i\geq 0$ and for any monomial $\eta$ of the SOHS part of $f$. Then
  1. (1)
    It is always possible to obtain a SOHS $\widetilde{f}$ by adding at most two nonconstant terms and a constant term with $f$ such that there exists a positive semidefinite Gram-like matrix of $\widetilde{f}$ which contains a positive semidefinite Gram-like matrix of the SOHS part of $f$ as a leading principal submatrix.
  2. (2)
    In fact, infinitely many such $\widetilde{f}$ exist.
Proof.
Let ${\rm deg}(f)=d$. By hypothesis, $f$ can be written as follows:
$f=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+a\zeta,$
where $a\in\mathbb{R}$, $g_{i}$s and $\zeta$ lie in $\mathbb{R}\langle\underline{X}\rangle$, ${\rm deg}(g_{i})\leq\tfrac{d}{2}$ if $d$ is even or ${\rm deg}(g_{i})\leq\tfrac{d-1}{2}$ if $d$ is odd, for all $i$ and $1\leq{\rm deg}(\zeta)\leq d$. Let $G$, say of order $k$, be a positive semidefinite Gram-like matrix for the SOHS part $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$ of $f$ such that there exist a column vector $W_{k}$ satisfying
  1. (a)
    $1$ does not appear as an entry of $W_{k}$,
  2. (b)
    $\zeta$ does not appear as an entry of $W_{k}$. and
  3. (c)
    $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}=W_{k}^{\ast}\;G\;W_{k}.$ (19)
It can be noted that existence of such $W_{k}$ satisfying
  • (a) is guaranteed by the assumption that $f$, hence $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$, doesn’t contain any constant term,
  • (b) is guaranteed by the hypothesis ${\rm rc}(\eta,i)\neq\zeta$ for any $i\geq 0$ for any monomial $\eta$ of $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$ (Indeed, let $\zeta$ appears as an entry of $W_{k}$. Then $\zeta^{\ast}\zeta$ appears in the expansion of $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$ as its coefficient in nonzero by (9b) of Theorem 3.5. But then ${\rm rc}(\zeta^{*}\zeta,{\rm deg}(\zeta))=\zeta$.),
  • (c) is guaranteed by Theorem 3.5.
Let us choose any $b,c\in\mathbb{R}$ such that
$\begin{split}b+c&>0,\text{\;and\;}\\ bc-a^{2}&>0.\end{split}$ (20)
Now we consider the matrix $\widetilde{G}$ given as follows:
$\widetilde{G}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}b&a\\ a&c\end{matrix}\end{array}\right)\;.$ (21)
We also define $\widetilde{W_{k+2}}$ as follows.
$\widetilde{W_{k+2}}:=\left(\begin{array}[]{@{}c}\begin{matrix}W_{k}^{t}&1&\zeta\end{matrix}\end{array}\right)^{t}\;.$ (22)
Then we have
$\widetilde{f}=f+a\;\zeta^{\ast}+c\;\zeta^{\ast}\zeta+b=\widetilde{W_{k+2}}^{\ast}\;\widetilde{G}\;\widetilde{W_{k+2}},$ (23)
where $\widetilde{G}$ and $\widetilde{W_{k+2}}$ are as in (21) and (22) respectively. Now as the matrices $G$, as in (19), and $\begin{pmatrix}b&a\\ a&c\\ \end{pmatrix}$, due to (20), are positive semidefinite matrices, so is the matrix $\widetilde{G}$ by Lemma 2.7. Hence $\widetilde{f}$, as in (23), is a SOHS by Theorem 3.5 and clearly obtained from $f$ by adding two nonconstant monomials $a\;\zeta^{\ast}$ and $c\;\zeta^{\ast}\zeta$ and a constant term $b$. Moreover if $\zeta^{\ast}=\zeta$, then at least the term $a\zeta^{\ast}$ is not needed in $\widetilde{f}$ In fact, in that case we can take $\widetilde{G}$ as
$\widetilde{G}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}b&\tfrac{a}{2}\\ \tfrac{a}{2}&c\end{matrix}\end{array}\right)\;,$
where $b+c>0$ and $04bc-a^{2}>$ 0.
As there are infinitely many choices for $b$ and $c$ satisfying (20), we have infinitely many choices for $\widetilde{G}$ and hence for $\widetilde{f}$. ∎
We now extend Theorem 3.15 when non-SOHS part of $f$ is having multiple terms.
Theorem 3.16.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial, without any constant term, which can be expressed as a SOHS plus $r$ many terms $a_{1}\zeta_{1},a_{2}\zeta_{2},\ldots,a_{r}\zeta_{r}$. Assume that ${\rm rc}(\eta,i)\neq\zeta_{j}$, $1\leq j\leq r$, for any $i\geq 0$ and for any monomial $\eta$ of the SOHS part of $f$. Then
  1. (1)
    It is always possible to obtain a SOHS $\widetilde{f}$ by adding at most $12r+$ 1 terms with $f$ such that there exists a positive semidefinite Gram-like matrix of $\widetilde{f}$ which contains a positive semidefinite Gram-like matrix of the SOHS part of $f$ as a leading principal submatrix.
  2. (2)
    In fact, infinitely many such $\widetilde{f}$ can be obtained.
Proof.
Let ${\rm deg}(f)=d$. By hypothesis, $f$ can be written as follows:
$f=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+\sum_{j=1}^{r}a_{j}\zeta_{j},$
where $a\in\mathbb{R}$, $g_{i}$s and $\zeta$ lie in $\mathbb{R}\langle\underline{X}\rangle$, ${\rm deg}(g_{i})\leq\tfrac{d}{2}$ if $d$ is even or ${\rm deg}(g_{i})\leq\tfrac{d-1}{2}$ if $d$ is odd, for all $1\leq i\leq m$ and $1\leq{\rm deg}(\zeta_{j})\leq d$ for all $1\leq j\leq r$. Let $G$, say of order $k$, be a positive semidefinite Gram-like matrix for the SOHS part $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$ of $f$ such that there exist a column vector $W_{k}$ satisfying
  1. (a)
    $1$ does not appear as an entry of $W_{k}$,
  2. (b)
    $\zeta_{1},\ldots,\zeta_{r}$ do not appear as entries of $W_{k}$, and
  3. (c)
    $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}=W_{k}^{\ast}\;G\;W_{k}.$ (24)
It can be noted that existence of such $W_{k}$ satisfying
  • (a) is guaranteed by the assumption that $f$, hence $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$, doesn’t contain any constant term,
  • (b) is guaranteed by the hypothesis ${\rm rc}(\eta,i)\neq\zeta_{j}$, $1\leq j\leq r$, for any $i\geq 0$ and for any monomial $\eta$ of the SOHS part of $f$ (Indeed, let $\zeta_{j}$ appears as an entry of $W_{k}$ for some $j$. Then $\zeta_{j}^{\ast}\zeta_{j}$ appears in the expansion of $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$ as its coefficient in nonzero by (9b) of Theorem 3.5. But then ${\rm rc}(\zeta_{j}^{\ast}\zeta_{j},{\rm deg}(\zeta_{j}))=\zeta_{j}$.),
  • (c) is guaranteed by Theorem 3.5.
Now we consider the matrix $\widetilde{G}$ given as follows:
$\widetilde{G}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}\sum_{j=1}^{r}a_{i}^{2}+1&a_{1}&a_{2}&\cdots&a_{r}\\ a_{1}&1&0&\cdots&0\\ a_{2}&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{r}&0&0&\cdots&1\end{matrix}\end{array}\right)\;.$ (25)
The leading principal minors $M_{l}$ of size $l$, $11\leq l\leq r+$ 1, of the submatrix
$\begin{pmatrix}\sum_{j=1}^{r}a_{i}^{2}+1&a_{1}&a_{2}&\cdots&a_{l-1}\\ a_{1}&1&0&\cdots&0\\ a_{2}&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{l-1}&0&0&\cdots&1\end{pmatrix}$ (26)
of the matrix $\widetilde{G}$, as defined in (25), are given as follows:
$M_{l}=\left\{\begin{array}[]{ll}\sum_{j=1}^{r}a_{i}^{2}+1&\mbox{for $l=1$};\\ \\ \sum_{j=l}^{r}a_{i}^{2}+1&\mbox{for $2\leq l\leq r$};\\ \\ 1&\mbox{for $l=r+1$}.\end{array}\right.$
Therefore the submatrix, as in (26), is positive definite by Theorem 2.4. Moreover, as the matrix $G$, as in (24), is positive semidefinite, so is the matrix $\widetilde{G}$ by Lemma 2.7.
We now define $\widetilde{W_{k+r+1}}$ as follows.
$\widetilde{W_{k+r+1}}:=\left(\begin{array}[]{@{}c}\begin{matrix}W_{k}^{t}&1&\zeta_{1}&\cdots&\zeta_{r}\end{matrix}\end{array}\right)^{t}\;.$ (27)
We define $\widetilde{f}$ as follows :
$\widetilde{f}:=\widetilde{W_{k+r+1}}^{\ast}\;\widetilde{G}\;\widetilde{W_{k+r+1}},$ (28)
where $\widetilde{G}$ and $\widetilde{W}_{k+r+1}$ are as in (25) and (27) respectively. Then we have
$\begin{split}\widetilde{f}&=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+\sum_{j=1}^{r}a_{j}\zeta_{j}+\sum_{j=1}^{r}a_{j}\zeta_{j}^{\ast}+\sum_{j=1}^{r}\zeta_{j}^{\ast}\zeta_{j}+\sum_{j=1}^{r}a_{i}^{2}+1\\ &=f+\sum_{j=1}^{r}a_{j}\zeta_{j}^{\ast}+\sum_{j=1}^{r}\zeta_{j}^{\ast}\zeta_{j}+\sum_{j=1}^{r}a_{i}^{2}+1.\end{split}$ (29)
It can be noted that $\widetilde{f}$ is obtained by adding $12r+$ 1 many terms with $f$. Moreover, if $\zeta_{i}$ is symmetric for any $i$, then we get the same $\widetilde{f}$ by adding even lesser number of terms. Therefore, first part of the assertion now follows from Theorem 3.5, (28) and (29).
It can be observed that any $\widetilde{G}$ obtained by replacing the $(k+1,k+1)$-th entry of the matrix by any $A>\sum_{j=1}^{r}a_{i}^{2}$ will produce another $\widetilde{f}$ satisfying the first part of the assertion. Hence, we have infinitely many choices for such $\widetilde{f}$. ∎
Remark 3.17.
  1. (1)
    It is clear from the construction of $\widetilde{G}$, as in (25), that even if we start with some other positive semidefinite Gram-like matrix $G^{\prime}$ of the SOHS part of $f$, though we get some different positive semidefinite Gram-like matrix of $\widetilde{f}$ but the polynomial $\widetilde{f}$ will still be the same.
  2. (2)
    The matrix $\widetilde{G}$, as in (25), could also have been constructed as follows:
    $\widetilde{G}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}A_{1}&a_{1}&a_{2}&\cdots&a_{r}\\ a_{1}&A_{2}&0&\cdots&0\\ a_{2}&0&A_{3}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{r}&0&0&\cdots&A_{r+1}\end{matrix}\end{array}\right)\;,$
    where $\prod_{i=1}^{r+1}A_{i}\geq\sum_{i=1}^{r}a_{i}^{2}$. In this case, though the newly added block is not positive definite in general but it is positive semidefinite.
The next theorem is towards getting an SOHS extension of $f$ not only with all the expected properties as before but also of arbitrarily high degree.
Theorem 3.18.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be as in Theorem 3.16. Then it is always possible to obtain a SOHS $\tilde{f}\in\mathbb{R}\langle\underline{X}\cup X_{n+1}\rangle$ of any possible even degree $\geq{\rm deg}(f)$, by adding at most $22r+$ 2 many terms with $f$, such that there exists a positive semidefinite Gram-like matrix of $\widetilde{f}$ which contains a positive semidefinite Gram-like matrix of $f$ as a leading principal submatrix.
Proof.
Let us keep the notations as in Theorem 3.16. We consider the matrix $\widetilde{G}$ given as follows:
$\widetilde{G}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}\sum_{j=1}^{r}a_{i}^{2}+1&a_{1}&a_{2}&\cdots&a_{r}&0\\ a_{1}&1&0&\cdots&0&0\\ a_{2}&0&1&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ a_{r}&0&0&\cdots&1&0\\ 0&0&0&\cdots&0&1\end{matrix}\end{array}\right)\;.$ (30)
The leading principal minors $M^{\prime}_{l}$ of size $l$, $21\leq l\leq r+$ 2, of the submatrix
$\begin{pmatrix}\sum_{j=1}^{r}a_{i}^{2}+1&a_{1}&a_{2}&\cdots&a_{l-2}&a_{l-1}\\ a_{1}&1&0&\cdots&0&0\\ a_{2}&0&1&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ a_{l-2}&0&0&\cdots&1&0\\ a_{l-1}&0&0&\cdots&0&1\end{pmatrix}$ (31)
of the matrix $\widetilde{G}$, as defined in (30), are given as follows:
$M^{\prime}_{l}=\left\{\begin{array}[]{ll}\sum_{j=1}^{r}a_{i}^{2}+1&\mbox{for $l=1$};\\ \\ \sum_{j=l}^{r}a_{i}^{2}+1&\mbox{for $2\leq l\leq r$};\\ \\ 1&\mbox{for $l=r+1$};\\ \\ 1&\mbox{for $l=r+2$ (as $a_{r+1}=0$)}.\end{array}\right.$
Therefore, the submatrix, as in (31), is positive definite by Theorem 2.4. Moreover, as the matrix $G$ is positive semidefinite, so is the matrix $\widetilde{G}$ by Lemma 2.7.
Let $p$ be any positive integer such that $2p\geq{\rm deg}(f)$. We also define $\widetilde{W_{k+r+2}}$ as follows.
$\widetilde{W_{k+r+2}}:=\left(\begin{array}[]{@{}c}\begin{matrix}W_{k}^{t}&1&\zeta_{1}&\cdots&\zeta_{r}&X_{n+1}^{p}\end{matrix}\end{array}\right)^{t}\;.$ (32)
Then the SOHS extension $\widetilde{f}$ of $f$, defined as
$\widetilde{f}:=\widetilde{W_{k+r+2}}^{\ast}\;\widetilde{G}\;\widetilde{W_{k+r+2}},$
where $\widetilde{G}$ and $\widetilde{W_{k+r+2}}$ are as in (30) and (32) respectively, satisfies the assertion. ∎
3.3. Positive semidefinite completion problem and SOHS via graphs
Positive semidefinite completion problem is one of the most important problems not only in the realm of mathematics but also in physics. Let $A$ be a $n\times n$ partially specified positive semidefinite matrix, i.e, a matrix with some unknown entries and some natural conditions on the known entries. Positive semidefinite completion problem deals with finding the pattern of known entries (and hence of the unknown entries as well) for which $A$ can be completed to a positive semidefinite matrix. Apart from having huge applications, this problem is very intriguing on its own. For example, a solution of this problem has been given via graph theory. It is known that such a solution exists when the pattern of known entries is given by a chordal graph.
In this subsection, we answer Question 1 via positive semidefinite completion problem. We recall some basic definitions and results related to this and study the same problem from the context of extending a partial SOHS polynomial to a SOHS polynomial.
Definition 3.19.
A partially specified $n\times n$ matrix is a partial symmetric matrix if
  • all the diagonal entries are specified, and
  • if the $(i,j)$-th entry is specified, then so is the $(j,i)$-th entry and they are equal.
Definition 3.20.
A partial symmetric matrix is said to be a partial positive definite matrix (respectively partial positive semidefinite matrix) if any of its completely specified principal minor is positive (respectively nonnegative).
Definition 3.21.
The specification graph $\mathcal{G}_{A}$ of a $n\times n$ partial symmetric matrix $A$ is a graph with vertex set $V=\{1,2,\ldots,n\}$ and for $i,j\in V$ with $i\neq j$, the vertices $i$ and $j$ are adjacent whenever $(i,j)$-th entry of $A$ is specified.
Definition 3.22.
A graph $\mathcal{G}$ with vertex set $\{1,2,\ldots,n\}$ is positive definite completable (respectively positive semidefinite completable) if any $n\times n$ partial positive definite (respectively semidefinite) matrix with specification graph $\mathcal{G}$ is completable to a positive definite (respectively semidefinite) matrix.
Definition 3.23.
A graph is said to be chordal if it doesn’t contain the cycle $\mathcal{C}_{k}$ on $k$ vertices as an induced subgraph for any $k\geq 4$.
Example 3.24.
  1. (1)
    The complete graph $\mathcal{K}_{n}$ on $n$ vertices, trees are some examples of chordal graphs.
  2. (2)
    The graph $\mathcal{C}_{4}$, the cycle on $4$ vertices, is not a chordal graph.
Theorem 3.25.
A graph $\mathcal{G}$ is positive definite (semidefinite) completable if and only if $\mathcal{G}$ is chordal.
Proof.
See [15, Theorem 7, p. 120] and [3, Lemma 12.7, p. 161 and Theorem 12.11, p. 163]. ∎
Now we describe what one means by incompleteness of a symmetric noncommutative polynomial w.r.t its partially specified Gram matrix and its completion.
Definition 3.26.
  1. (1)
    A symmetric noncommutative polynomial $f$ of degree $2d$ is said to be incomplete w.r.t its representation $\mathbf{R}_{f}$, as in (1), along with conditions as in (2), if
    • all $a_{i}$s are specified, and
    • some $b_{ij}$s are unspecified.
  2. (2)
    A symmetric noncommutative polynomial $f$ of degree $2d$ is said to be incomplete w.r.t its representation $\mathbf{R}_{f}$, as in (10), along with conditions as in (11), if
    • all $a_{i}$s are specified (positive real numbers by conditions as in (11b)), and
    • some $b_{ij}$s are unspecified.
Definition 3.27.
  1. (1)
    Let $W_{s(d,n)}$ be a column vector of size $s(d,n)$ of all monomials from $\mathbb{R}\langle\underline{X}\rangle$ of degree at most $d$ taken in lexicographic order and $G$ a $s(d,n)\times s(d,n)$ partial positive semidefinite matrix. Then for any positive semidefinite completion $\overline{G}$ of $G$, the polynomial $W_{s(d,n)}^{\ast}\;\overline{G}\;W_{s(d,n)}$ is called a SOHS completion of the polynomial $f:=W_{s(d,n)}^{\ast}\;G\;W_{s(d,n)}$ w.r.t partial positive semidefinite Gram matrix $G$, and is denoted by $\overline{f}^{\;G}$.
  2. (2)
    Let $W_{k}$ be a column vector of size $k$ of all monomials from $\mathbb{R}\langle\underline{X}\rangle$ of degree at most $d$ taken in any arbitrary order and $G$ a $k\times k$ partial positive semidefinite matrix with positive diagonal entries. Then for any positive semidefinite completion $\overline{G}$ of $G$, the polynomial $W_{k}^{\ast}\;\overline{G}\;W_{k}$ is called a SOHS completion of the polynomial $f:=W_{s(d,n)}^{\ast}\;G\;W_{s(d,n)}$ w.r.t partial positive semidefinite Gram-like matrix $G$, and is denoted by $\overline{f}^{\;G,\;W_{k}}$.
We now want to look at incomplete polynomials that are not SOHS, but potentially very close to SOHS polynomials. That is, whenever we collect appropriate monomials (with known coefficients) of such an incomplete polynomial $f$ which have the potential to form a SOHS, it does so.
Definition 3.28.
  1. (1)
    Let $f$ be a incomplete polynomial w.r.t a given representation $\mathbf{R}_{f}$, as in (1), along with conditions as in (2). Then $f$ is said to be quasi SOHS w.r.t $\mathbf{R}_{f}$ if for any $K\subset\{1,2,\ldots,s(d,n)\}$ such that $b_{ij}$s are known for all $i,j\in K$,
    $\sum_{i\in K}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{i,j\in K,i<j}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})$
    is a SOHS.
  2. (2)
    Let $f$ be a incomplete polynomial w.r.t a given representation $\mathbf{R}_{f}$, as in (10), along with conditions as in (11). Then $f$ is said to be quasi SOHS w.r.t $\mathbf{R}_{f}$ if for any $K\subset\{1,2,\ldots,k\}$ such that $b_{ij}$s are known for all $i,j\in K$,
    $\sum_{i\in K}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{i,j\in K,i<j}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})$
    is a SOHS.
We are now in a stage to reformulate Theorem 3.25 in the context of the dictum of this paper. To be specific, we provide a criteria regarding when a quasi SOHS polynomial qualifies to a SOHS polynomial. We have the following theorem in that regard.
Theorem 3.29.
  1. (1)
    Let $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ be a quasi SOHS polynomial w.r.t a representation $\mathbf{R}_{f}$, as in (1), along with conditions as in (2). Then any such $f$ has a SOHS completion $\overline{f}^{\;G_{f}}$ w.r.t the partial positive semidefinite Gram matrix $G_{f}$ associated with $\mathbf{R}_{f}$ if and only if the specification graph of $G_{f}$ is chordal.
  2. (2)
    Let $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ be a quasi SOHS polynomial w.r.t a representation $\mathbf{R}_{f}$, as in (10), along with conditions as in (11). Then any such $f$ has a SOHS completion $\overline{f}^{\;G_{f},\;W_{k}}$ w.r.t the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$ if and only if the specification graph of $G_{f}$ is chordal. Here $W_{k}$ is the column vector $\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}\end{pmatrix}^{t}$, $\zeta_{i}$, $1\leq i\leq k$, being the monomials appearing in $\mathbf{R}_{f}$.
Proof.
(1)   Let $f$ be a quasi SOHS polynomial of degree $2d$ w.r.t a representation
$\mathbf{R}_{f}\;\;:\;\;f=\sum_{i=1}^{s(d,n)}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq s(d,n)}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$
where $a_{i},b_{ij}\in\mathbb{R}$, $\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle$ are all monomials of degree at most $d$. Therefore, for any $K\subset\{1,2,\ldots,s(d,n)\}$ such that $b_{ij}$s are known for all $i,j\in K$,
$\sum_{i\in K}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{i,j\in K,i<j}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})$
is a SOHS. Hence, for any such set $K=\{i_{0},i_{1},\ldots,i_{\mid K\mid}\}$ with $i_{0}<i_{1}<\cdots<i_{\mid K\mid}$, the matrix
$\begin{pmatrix}a_{i_{1}}&b_{i_{1}i_{2}}&\cdots&b_{i_{1}i_{\mid K\mid}}\\ b_{i_{1}i_{2}}&a_{i_{2}}&\cdots&b_{i_{2}i_{\mid K\mid}}\\ \vdots&\vdots&\ddots&\vdots\\ b_{i_{1}i_{\mid K\mid}}&b_{i_{2}i_{\mid K\mid}}&\cdots&a_{i_{\mid K\mid}}\end{pmatrix}$
is positive semidefinite by Theorem 3.5. Therefore any specified principal minor of the Gram matrix $G_{f}$, as in (4), associated with $\mathbf{R}_{f}$ is positive semidefinite, that is to say $G_{f}$ is a partial positive semidefinite matrix. Therefore, it makes sense to talk about SOHS completion of $f$ w.r.t $G_{f}$ following Definition 3.27. Now, the assertion follows from Theorem 3.25.
(2)   Follows similarly. ∎
Remark 3.30.
  1. (1)
    If we take the specification graph of $G_{f}$ as the complete graphs $\mathcal{K}_{s(d,n)}$ and $\mathcal{K}_{k}$, then we get back Theorem 2.12 and Theorem 3.5 respectively. That is to say, Theorem 3.29 generalises Theorem 2.12 and Theorem 3.5.
  2. (2)
    Let $\mathcal{G}$ be nonchordal graph. Then also there might exist a quasi SOHS polynomial $f$ with $\mathcal{G}$ as the specification graph of its Gram matrix $G_{f}$ which has a SOHS completion. For example, let $\zeta_{1}=1$, $\zeta_{2}=X_{1}X_{2}$, $\zeta_{3}=X_{2}X_{1}$ and $\zeta_{4}=X_{2}^{2}$ and $W_{4}=\begin{pmatrix}1&X_{1}X_{2}&X_{2}X_{1}&X_{2}^{2}\end{pmatrix}^{t}$. Consider the following polynomial $f$ given as
    $\begin{split}f=&5\zeta_{1}^{\ast}\zeta_{1}+5\zeta_{2}^{\ast}\zeta_{2}+5\zeta_{3}^{\ast}\zeta_{3}++5\zeta_{4}^{\ast}\zeta_{4}+5(\zeta_{1}^{\ast}\zeta_{2}+\zeta_{2}^{\ast}\zeta_{1})+\ast(\zeta_{1}^{\ast}\zeta_{3}+\zeta_{3}^{\ast}\zeta_{1})\\ &+5(\zeta_{1}^{\ast}\zeta_{4}+\zeta_{4}^{\ast}\zeta_{1})+5(\zeta_{2}^{\ast}\zeta_{3}+\zeta_{3}^{\ast}\zeta_{2})+\ast(\zeta_{2}^{\ast}\zeta_{4}+\zeta_{4}^{\ast}\zeta_{2})+5(\zeta_{3}^{\ast}\zeta_{4}+\zeta_{4}^{\ast}\zeta_{3})\\ =5&+(5+\ast)X_{1}X_{2}+(5+\ast)X_{2}X_{1}+10X_{2}^{2}+5X_{1}X_{2}^{2}X_{1}+5X_{2}X_{1}^{2}X_{2}+5X_{1}X_{2}X_{1}X_{2}\\ &+5X_{2}X_{1}X_{2}X_{1}+5X_{1}X_{2}^{3}+5X_{2}^{3}X_{1}+\ast X_{2}X_{1}X_{2}^{2}+\ast X_{2}^{2}X_{1}X_{2}+5X_{2}^{4}.\end{split}$ (33)
    It is easy to check that $f$ is a quasi SOHS polynomial of degree $4$ w.r.t the given representation and the matrix $G_{f}$ given by
    $G_{f}=\begin{pmatrix}5&5&\ast&5\\ 5&5&5&\ast\\ \ast&5&5&5\\ 5&\ast&5&5\\ \end{pmatrix}$
    is a partial positive semidefinite Gram-like matrix of $f$. The matrix $G_{f}$ is also a partial positive semidefinite matrix as $f$ is quasi SOHS and by Theorem 3.5. It can also be noted that the specification graph of $G_{f}$ is $\mathcal{C}_{4}$, which is not chordal. Still there exists a SOHS completion $\overline{f}^{\;G_{f},\;W_{4}}$, which is obtained by taking =5$\ast$=5∗ = 5 in (33) as the positive semidefinite matrix
    $\overline{G_{f}}=\begin{pmatrix}5&5&5&5\\ 5&5&5&5\\ 5&5&5&5\\ 5&5&5&5\\ \end{pmatrix}$
    is clearly a positive semidefinite completion of $G_{f}$. But, this doesn’t contradict Theorem 3.29.
    To see this, we first note that determinant of the matrix
    $\begin{pmatrix}5&5&\ast\\ 5&5&5\\ \ast&5&5\end{pmatrix}$ (34)
    is $-(\ast-5)^{2}$. Therefore, the matrix as in (34) is positive semidefinite if and only if =5$\ast$=5∗ = 5. That is to say, the only positive semidefinite completion of this partial positive semidefinite matrix is
    $\begin{pmatrix}5&5&5\\ 5&5&5\\ 5&5&5\end{pmatrix}\;.$
    We now exploit this uniqueness. Consider the polynomial
    $\begin{split}f^{\prime}=&5\zeta_{1}^{\ast}\zeta_{1}+5\zeta_{2}^{\ast}\zeta_{2}+5\zeta_{3}^{\ast}\zeta_{3}+5\zeta_{4}^{\ast}\zeta_{4}+5(\zeta_{1}^{\ast}\zeta_{2}+\zeta_{2}^{\ast}\zeta_{1})+\ast(\zeta_{1}^{\ast}\zeta_{3}+\zeta_{3}^{\ast}\zeta_{1})\\ &+\sqrt{5}(\zeta_{1}^{\ast}\zeta_{4}+\zeta_{4}^{\ast}\zeta_{1})+5(\zeta_{2}^{\ast}\zeta_{3}+\zeta_{3}^{\ast}\zeta_{2})+\ast(\zeta_{2}^{\ast}\zeta_{4}+\zeta_{4}^{\ast}\zeta_{2})+5(\zeta_{3}^{\ast}\zeta_{4}+\zeta_{4}^{\ast}\zeta_{3})\\ =5&+(5+\ast)X_{1}X_{2}+(5+\ast)X_{2}X_{1}+2\sqrt{5}X_{2}^{2}+5X_{1}X_{2}^{2}X_{1}+5X_{2}X_{1}^{2}X_{2}+5X_{1}X_{2}X_{1}X_{2}\\ &+5X_{2}X_{1}X_{2}X_{1}+5X_{1}X_{2}^{3}+5X_{2}^{3}X_{1}+\ast X_{2}X_{1}X_{2}^{2}+\ast X_{2}^{2}X_{1}X_{2}+5X_{2}^{4}.\end{split}$ (35)
    Again it is a qusi SOHS polynomial of degree $4$ and associated partial positive semidefinite Gram-like matrix $G_{f^{\prime}}$ is given by
    $G_{f^{\prime}}=\begin{pmatrix}5&5&\ast&\sqrt{5}\\ 5&5&5&\ast\\ \ast&5&5&5\\ \sqrt{5}&\ast&5&5\\ \end{pmatrix}$ (36)
    having specification graph $\mathcal{C}_{4}$. Suppose, the matrix $G_{f^{\prime}}$ is positive semidefinite completable. Then first consider the leading principal submatrix of $G_{f^{\prime}}$ obtained by deleting last row and column. Then, as this has to be positive semidefinite after completion, $\ast$ in $(1,3)$-th and $(3,1)$-th position must have to be $5$ by the aforementioned uniqueness. Now, considering the principal submatrix obtained by deleting first row and column and applying same argument, we obtain $\ast$ in $(2,4)$-th and $(4,2)$-th position must have to be $5$. But then the principal submatrix of $G_{f^{\prime}}$, obtained either by deleting second row and second column (or by deleting third row and third column), has determinant $(\sqrt{5}-5)(25-5\sqrt{5})(<0)$. Therefore, the matrix $G_{f^{\prime}}$ is not positive semidefinite completable and hence the polynomial, as in (35) doesn’t have a SOHS completion.
We now go through a couple of applications of Theorem 3.29.
Proposition 3.31.
Let $f$ be a quasi SOHS polynomial of degree $2d$ w.r.t to the representation $\mathbf{R}_{f}$
$\mathbf{R}_{f}\;:\;f=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$
as in (10) along with conditions as in (11). If out of $\tfrac{k(k-1)}{2}$ many $b_{ij}$s, any $k-1$ are specified and remaining $\tfrac{(k-1)(k-2)}{2}$ are unspecified, and if there is no column (or row) of the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$ with diagonal entry as the only specified entry, then $f$ has a SOHS completion.
Proof.
By hypothesis, in any column of $G_{f}$ other than the diagonal entry at least one entry is specified. This ensures that the specification graph $\mathcal{G}_{G_{f}}$ of $G_{f}$ is connected. Moreover, the hypothesis that any $k-1$ many $b_{ij}$s are specified affirms that there are exactly $k-1$ edges between $k$ many vertices of $\mathcal{G}_{G_{f}}$. Therefore, the graph $\mathcal{G}_{G_{f}}$ is indeed a tree and hence chordal. Hence, the assertion follows from Theorem 3.29. ∎
Corollary 3.32.
Let $f$ be a quasi SOHS polynomial of degree $2d$ w.r.t to the representation $\mathbf{R}_{f}$
$\mathbf{R}_{f}\;:\;f=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$
as in (10) along with conditions as in (11). If out of $\tfrac{k(k-1)}{2}$ many $b_{ij}$s, exactly $k-1$ are specified for some fixed $i$ (or fixed $j$), then $f$ has a SOHS completion.
Proof.
Follows from Proposition 3.31. ∎
Remark 3.33.
It can be noted that, taking $k=4$ in Corollary 3.32, the specification graphs $\mathcal{G}_{G_{f}}$ are the following rooted trees of height one on $4$ vertices.
[…]
3.4. Positive semidefinite completion problem and SOHS via algebraic sets
In Subsection 3.3, we answered Question 1 using techniques of solving positive semidefinite completion problem via graph theory. Recall that, given a partial positive semidefinite matrix $G$, we considered a graph associated with it which was obtained by keeping track of the specified off diagonal entries of $G$. On the contrary, now we answer Question 1 using another method of solving positive semidefinite completion problem. We do so by associating a projective algebraic set to it that keeps track of the unspecified off diagonal entries. In the process, we obtain a nice connection between noncommutative polynomial and usual (commutative) polynomial.
In this section, by a polynomial we mean the usual commutative polynomial. We denote commutative variables by small alphabets, unlike the noncommutative variables. By a form we mean a homogeneous polynomial. It is obvious that if a (homogeneous) polynomial has a sum of square (SOS) representation then it should be nonnegative. But the converse is not true. In fact, Hilbert, in 1888, proved that converse is true only for quadratic forms, binary forms and ternary quartics (cf. [19]). Many years later, Motzkin gave a specific example of ternary sextic, namely $x_{0}^{6}+x_{1}^{4}x_{2}^{2}+x_{1}^{2}x_{2}^{4}-3x_{0}^{2}x_{1}^{2}x_{2}^{2}$, which is a nonnegative polynomial but can not be expressed as a SOS. In 2016, Blekherman, Smith and Velasco has generalised Hilbert’s result. Instead of looking into globally nonnegative polynomials, they looked at nonnegative quadratic forms over a nondegenerate totally-real projective variety $X\subset\mathbb{P}^{n}$. Here, by a totally-real variety they meant that the set of real points in $X$ is a dense subset of set of complex points of $X$. They proved that any nonnegative polynomial would be a sum of squares, modulo the ideal of $X$ if and only if $X$ is a variety of minimal degree, i.e, ${\rm deg}(X)={\rm codim}(X)+1$ (cf. [8, Theorem 1.1,p. 893]). In 2017, Blekherman, Sinn and Velasco generalised it even further for any nondegenerate totally-real projective algebraic set. As the notion of degree doesn’t make sense for reducible algebraic sets, they observed that a right generalisation of the notion of degree is Castelnuovo-Mumford regularity. We now recall the meaning of this notion for a projective algebraic set.
Definition 3.34.
Let $k$ be a field of characteristic 00. A finitely generated $k[x_{0},x_{1},\ldots,x_{n}]$-module $M$ has a unique free resolution
$\cdots\rightarrow F_{t}\rightarrow F_{t-1}\rightarrow\cdots\rightarrow F_{0}\rightarrow M\rightarrow 0.$
Let $b_{j}$ be the largest degree of a minimal generator in the module $F_{j}$. We also fix the convention that $b_{j}=-\infty$ if $F_{j}$ is the zero module. Then the module $M$ is said to be $m$-regular if $b_{j}\leq m+j$ for all $j$.
Definition 3.35.
A projective algebraic set $X\subset\mathbb{P}^{n}$ is said to be $m$-regular if the defining ideal $I$ of $X$ is $m$-regular as a $k[x_{0},x_{1},\ldots,x_{n}]$-module.
We now state the main result obtained by Blekherman, Sinn and Velasco generalising [8, Theorem 1.1,p. 893].
Theorem 3.36.
Let $X\subset\mathbb{P}^{n}$ be a totally-real projective algebraic set. Every nonnegative quadratic form on $X$ is a sum of squares, modulo the ideal of $X$ if and only if $X$ is 2-regular.
Proof.
See [6, Theorem 9, p. 181]. ∎
Remark 3.37.
It can be noted that linear projective varieties are $1$-regular. Apart from that, any varieties of minimal degree are $2$-regular. That is why, the apt generalisation of the notion of degree for a reduced algebraic set is regularity.
Apparently, it might seem that the results proved in [8] and [6] are only for quadratic forms, whereas Hilbert proved the same for any $2d$ degree polynomials. But, they showed that it is enough to consider quadratic forms as any other even degree forms can be settled down using Veronese embedding. To be precise, to look at degree $2d$ forms on a variety $X$, it is enough to consider quadratic forms on the image $\nu_{d}(X)$ of $X$ via $d$-th Veronese embedding $\nu_{d}$, (cf. [7, p. 737]).
Both of the results, i.e, Theorem 1.1 in [8] and Theorem 9 in [6], have a deep connection with positive semidefinite completion problem. Recall that, given a partial positive semidefinite matrix $G$, its specification graph encodes the known part of $G$. On the contrary, to encode the unknown part of $G$ a projective algebraic set can be associated with it. In this regard, we have the following definition.
Definition 3.38.
The subspace arrangement $V_{A}$ of a $n\times n$ partial symmetric matrix $A$ is a projective algebraic subset of $\mathbb{P}^{n-1}$ whose vanishing homogeneous ideal is generated by the (square free) quadratic monomials $x_{i-1}x_{j-1}$, $1\leq i<j\leq n$, whenever $(i,j)$-th entry of $A$ is unspecified.
As mentioned earlier, it is expected that given a partial symmetric matrix, its specification graph and subspace arrangement should be related. Following result by Fröberg points out that relation. In the process, it provides a bridge between algebraic geometry, commutative algebra, graph theory and positive semidefinite completion problem.
Theorem 3.39.
Let $G$ be a partial symmetric matrix. Then the subspace arrangement $V_{G}$ is $2$-regular if and only if the specification graph $\mathcal{G}_{G}$ of $G$ is chordal.
Proof.
See [13]. ∎
Unlike commutative setup, the relation between nonnegative noncommutative polynomials and SOHS polynomials is simple. In fact, Helton, in 2002, proved that any positive noncommutative polynomial is a SOHS polynomial (cf. [16]). But the equality between the nonnegative quadratic forms and SOS modulo a suitable projective algebraic set has the potential to allow a SOHS extension of a quasi SOHS polynomial of any even degree. We encode this in the following theorem, which is nothing but a reinterpretation of Theorem 3.29 through the lens of algebraic geometry and commutative algebra. Before that, we quickly recall a couple of definitions along with their notations.
We denote by $S$ the homogeneous coordinate ring of a projective algebraic set $X\subset\mathbb{P}^{n}$. That is, if $I$ is the homogeneous ideal of $\mathbb{R}[x_{0},\ldots,x_{n}]$ generated by all forms that vanish on $X$, then $S:=\tfrac{\mathbb{R}[x_{0},\ldots,x_{n}]}{I}$. By $S_{j}$ we denote the vector space over $\mathbb{R}$ spanned by the degree $j$ forms of $S$. Then by $P_{X}$ and $\Sigma_{X}$ we denote the following two sets.
[…]
Theorem 3.40.
  1. (1)
    Let $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ be a quasi SOHS polynomial w.r.t a representation $\mathbf{R}_{f}$, as in (1), along with conditions as in (2). Then the following are equivalent:
    1. (a)
      Any such $f$ has a SOHS completion $\overline{f}^{\;G_{f}}$ w.r.t the partial positive semidefinite Gram matrix $G_{f}$ associated with $\mathbf{R}_{f}$.
    2. (b)
      The subspace arrangement $V_{G_{f}}(\subseteq\mathbb{P}^{s(d,n)-1})$ of $G_{f}$ satisfies $P_{V_{G_{f}}}=\Sigma_{V_{G_{f}}}$.
    3. (c)
      The subspace arrangement $V_{G_{f}}(\subseteq\mathbb{P}^{s(d,n)-1})$ of $G_{f}$ is $2$-regular.
  2. (2)
    Let $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ be a quasi SOHS polynomial w.r.t a representation $\mathbf{R}_{f}$, as in (10), along with conditions as in (11). Then the following are equivalent:
    1. (a)
      Any such $f$ has a SOHS completion $\overline{f}^{\;G_{f},\;W_{k}}$ w.r.t the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$.
    2. (b)
      The subspace arrangement $V_{G_{f}}(\subseteq\mathbb{P}^{k-1})$ of $G_{f}$ satisfies $P_{V_{G_{f}}}=\Sigma_{V_{G_{f}}}$.
    3. (c)
      The subspace arrangement $V_{G_{f}}(\subseteq\mathbb{P}^{k-1})$ of $G_{f}$ is $2$-regular.
    Here $W_{k}$ is the column vector $\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}\end{pmatrix}^{t}$, $\zeta_{i}$, $1\leq i\leq k$, being the monomials appearing in $\mathbf{R}_{f}$.
Proof.
Follows from Theorem 3.29, Theorem 3.36 and Theorem 3.39. ∎
We now apply Theorem 3.29 and Theorem 3.40 to provide a couple of examples of some quasi SOHS polynomials that admit SOHS completions.
Example 3.41.
Consider any quasi SOHS polynomial $f$ whose partial positive semidefinite Gram-like matrix $G_{f}$ is given as follows :
$G_{f}=\begin{pmatrix}5&5&5&\ast\\ 5&5&5&\ast\\ 5&5&5&5\\ \ast&\ast&5&5\\ \end{pmatrix}.$
Then the corresponding subspace arrangement $V_{G_{f}}$ is the projective algebraic set sitting inside $\mathbb{P}^{3}$ whose defining ideal is the homogeneous ideal $I:=\langle x_{0}x_{3},x_{1}x_{3}\rangle$. Using Macaulay2 software (cf. [14]), we compute the free resolution and Betti table of $I$, considered as a $\mathbb{R}[x_{0},x_{1},x_{2},x_{3}]$-module, as follows :
Free resolution of $I$ : Denoting $\mathbb{R}[x_{0},x_{1},x_{2},x_{3}]$ by $R$, we have :
[…]
Betti table of $I$:
0   1total :2   12:2   $\begin{split}&0\;\;\;1\\ \text{total :}\;&2\;\;\;1\\ 2:\;&2\;\;\;1\end{split}$
Therefore, we obtain that the ideal $I$ is $2$-regular and so is the projective algebraic set $V_{G_{f}}$. Therefore, by Theorem 3.40, any such quasi SOHS polynomial $f$ admits a SOHS completion w.r.t $G_{f}$. This is also evident from Theorem 3.29 as the specification graph $\mathcal{G}_{G_{f}}$
[…]
is chordal.
Example 3.42.
Consider any quasi SOHS polynomial $f$ whose partial positive semidefinite Gram-like matrix $G_{f}$ has the following specification graph $\mathcal{G}_{f}$ :
[…]
Then the corresponding subspace arrangement $V_{G_{f}}$ is the projective algebraic subset of $\mathbb{P}^{4}$ whose defining ideal is the homogeneous ideal $I:=\langle x_{0}x_{2},x_{1}x_{2},x_{2}x_{3}\rangle$. Using Macaulay2 software (cf. [14]), we compute the free resolution and Betti table of $I$, considered as a $\mathbb{R}[x_{0},x_{1},x_{2},x_{3},x_{4}]$-module, as follows :
Free resolution of $I$ : Denoting $\mathbb{R}[x_{0},x_{1},x_{2},x_{3},x_{4}]$ by $R$, we have :
[…]
Betti table of $I$:
0   1   2total :3   3   12:3   3   $\begin{split}&0\;\;\;1\;\;\;2\\ \text{total :}\;&3\;\;\;3\;\;\;1\\ 2:\;&3\;\;\;3\;\;\;1\end{split}$
Therefore, we obtain that the ideal $I$ is $2$-regular and so is the projective algebraic set $V_{G_{f}}$. Therefore, by Theorem 3.40, any such quasi SOHS polynomial $f$ admits a SOHS completion w.r.t $G_{f}$. In fact, it also follows from Theorem 3.29 as the specification graph $\mathcal{G}_{f}$ is chordal.
We now apply Theorem 3.29 and Theorem 3.40 to provide a couple of examples of some quasi SOHS polynomials that do not admit SOHS completions.
Example 3.43.
Consider the partial positive semidefinite matrix, as in (36),
$G^{\prime}=\begin{pmatrix}5&5&\ast&\sqrt{5}\\ 5&5&5&\ast\\ \ast&5&5&5\\ \sqrt{5}&\ast&5&5\\ \end{pmatrix}.$
Then the corresponding subspace arrangement $V_{G^{\prime}}$ is the projective algebraic set sitting inside $\mathbb{P}^{3}$ whose defining ideal is the homogeneous ideal $J:=\langle x_{0}x_{2},x_{1}x_{3}\rangle$. Using Macaulay2 software (cf. [14]), we compute the free resolution and Betti table of $I$, considered as a $\mathbb{R}[x_{0},x_{1},x_{2},x_{3}]$-module, as follows :
Free resolution of $J$ : Denoting $\mathbb{R}[x_{0},x_{1},x_{2},x_{3}]$ by $R$, we have :
[…]
Betti table of $J$:
0   1total :2   12:2.3:.   $\begin{split}&0\;\;\;1\\ \text{total :}\;&2\;\;\;1\\ 2:\;&2\;\;\;\textbf{.}\\ 3:\;&\textbf{.}\;\;\;1\\ \end{split}$
Therefore, we obtain that the ideal $J$ is $3$-regular and so is the projective algebraic set $V_{G^{\prime}}$. Therefore, by Theorem 3.40, there exists a quasi SOHS polynomial that doesn’t admit a SOHS completion w.r.t $G^{\prime}$. In fact, the polynomial $f^{\prime}$, as given in (35), is one such. This is not all surprising. In fact, it follows from Theorem 3.29 as the specification graph $G^{\prime}$
[…]
is the cycle $\mathcal{C}_{4}$ on four vertices, which is not a chordal graph.
Example 3.44.
Consider any partial positive semidefinite matrix $G$ which has the following specification graph $\mathcal{G}$ :
[…]
Then the corresponding subspace arrangenment $V_{G}$ is the projective algebraic set sitting inside $\mathbb{P}^{4}$ whose defining homogeneous ideal $J$ is the ideal $\langle x_{0}x_{2},x_{1}x_{3},x_{1}x_{4},x_{2}x_{4}\rangle$. Using Macaulay2 software (cf. [14]), we compute the free resolution and Betti table of $J$, considered as a $\mathbb{R}[x_{0},x_{1},x_{2},x_{3},x_{4}]$-module, as follows :
Free resolution of $J$ : Denoting $\mathbb{R}[x_{0},x_{1},x_{2},x_{3},x_{4}]$ by $R$, we have :
[…]
Betti table of $J$:
0   1   2total :4   4   12:4   3.3:.   1   $\begin{split}&0\;\;\;1\;\;\;2\\ \text{total :}\;&4\;\;\;4\;\;\;1\\ 2:\;&4\;\;\;3\;\;\;\textbf{.}\\ 3:\;&\textbf{.}\;\;\;1\;\;\;1\\ \end{split}$
Therefore, we obtain that the ideal $J$ is $3$-regular and so is the projective algebraic set $V_{G}$. Therefore, by Theorem 3.40, there exists a quasi SOHS polynomial that doesn’t admit a SOHS completion w.r.t $G$. In fact, it also follows from Theorem 3.29 as the specification graph $\mathcal{G}$ contains $\mathcal{C}_{4}$ as induced subgraph.
Remark 3.45.
The flexibility of a Gram-like matrix over a Gram matrix is quite evident from these above examples. As we deal with Gram-like matrices, we could play around with graphs having any number of vertices. For a Gram matrix, we can only deal with graphs with $s(d,n)$ many vertices for some $d\geq 0$ and $n>0$.
Remark 3.46.
As a next step, it would be interesting to generate algorithms out of the criteria provided in Theorem 3.16, Theorem 3.29 and Theorem 3.40 and use that further to produce SDP corresponding to these extension problems.
Acknowledgements
Both the authors extend their gratitude to Prof. Sumesh K and Prof. Arijit Dey for their suggestions regarding the article. The authors acknowledge Rajib Sarkar and Sreenanda S B for their assistance regarding the use of Macaulay2 software. The authors also thank Agnihotri Roy for reexamining some tedious calculations. The first named author would like to thank Indian Institute of Technology Madras for financial support (Office order No.F.ARU/R10/IPDF/2024). The second named author would like to thank Indian Institute of Technology Madras and ICSR for financial support (Project number - SB22231267MAETWO008573).