### v2

# Gram-like matrix preserving extensions and completions of noncommutative polynomials

Abstract.
Given a positive noncommutative polynomial $f$, equivalently a sum of Hermitian squares (SOHS), there exists a positive semidefinite Gram matrix that encrypts all the structural essence of $f$. There are no available methods for extending a noncommutative polynomial to a SOHS keeping the Gram matrices unperturbed. As a remedy, we introduce an equally significant notion of Gram-like matrices and provide linear algebraic techniques to get the desired extensions. We further use positive semidefinite completion problem to get SOHS and provide criteria in terms of chordal graphs and 2-regular projective algebraic sets.
Keywords : Noncommutative polynomial, sum of Hermitian squares, chordal graph, Castelnuovo-Mumford regularity, Betti table, free resolution, matrix completion.
2020 Mathematics Subject Classification : 14P05, 14N25, 13D02, 05C50, 11E25, 90C90.
1. Introduction
Hilbert’s 17th problem states whether every multivariate polynomial that takes only nonnegative values over the reals can be represented as sum of squares (SOS) of polynomials. Hilbert, in 1888, proved that it is true only for quadratic forms, binary forms and ternary quatrics in the commutative case. In the noncommutative setup, Helton proved that any polynomial is positive if and only if it is a sum of Hermitian squares (SOHS). The SOS and SOHS techniques are heavily used for polynomial optimization and convex optimization, in particular, in semidefinite programming (SDP). Many of the SOHS techniques are taken from real algebraic geometry and inspired from functional analysis.
The SOHS techniques apply across many different disciplines including control theory, probability and statistics. Out of many directions, hierarchy techniques are still very active areas of research in quantum information theory (cf. [12] and [23]) and have been proven useful in discriminating the quantum states. A large portion of the theory for noncommutative polynomials is being developed by Helton, McCullough, Klep and their collaborators (cf. [10], [17], [18], [20] and [21]). The book “Optimization of Polynomials in Non-commuting Variables” (cf. [9]) by Burgdorf, Klep and Povh is a beautiful introduction to the subject. Pablo Parillo has used SOS techniques in different SDP to produce different control theory oriented applications (cf. [24]). The work of Blekherman and his collaborators on SOS techniques provide fruitful insight about the fundamental geometric properties of some projective algebraic sets (cf. [8], [6] and [7]). Motivated by the theory of noncommutative polynomials, noncommutative functions have been developed over noncommutative sets and many researchers are working on the subject (cf. [1] and [26]).
Our main contributions are threefold. First, we provide a linear algebraic construction to extend noncommutative polynomials to SOHS, by adding necessary terms, that ensures the Gram-like matrix of the SOHS part is a principal submatrix of the extension’s matrix (cf. Theorem 3.24). Second, we leverage chordal graph theory to characterize when a partially specified noncommutative polynomial can be completed to a SOHS, linking matrix completion to SOHS completion (cf. Theorem 3.35). Third, we connect these completions to projective algebraic sets, showing that their 2-regularity guarantees SOHS completions (cf. Theorem 3.47). These results blend linear algebra, graph theory, and algebraic geometry. One of the key aspects that the reader will take away from this paper is some techniques to produce ample SOHS. Moreover, these techniques will be fruitful in polynomial optimization.
We now go through the chronology of the paper in a bit more detail. In Section 2, we recall some prerequisites and discuss the notion of a Gram matrix & monomial vector associated to a symmetric polynomial & SOHS. We introduce the notion of a Gram-like matrix which reduces the size of Gram matrix and disregards the order of the constituent monomials in the monomial vector (cf. Remark 3.9 for a detailed comparison between a Gram matrix and a Gram-like matrix). Moreover, it keeps the soul of a Gram matrix intact.
The following question captures the central theme of this paper:
Question 1.
Given a noncommutative polynomial $f\in\mathbb{R}\langle\underline{X}\rangle$ that includes SOHS components, can we construct a SOHS polynomial $\widetilde{f}\in\mathbb{R}\langle\underline{X}\rangle$ that extends (or completes) $f$ such that the positive semidefinite Gram-like matrices of the SOHS components are principal submatrices of the positive semidefinite Gram-like matrix of $\tilde{f}$?
Theorem 3.24, Theorem 3.35 and Theorem 3.47 answer Question 1 in three different set up in three consecutive subsections.
  1. (1)
    In subsection 3.2, assuming that there exists some $\widetilde{f}$ satisfying the properties as in Question 1, we first note that the structure of a positive semidefinite Gram-like matrix $G_{\widetilde{f}}$ of such $\widetilde{f}$ would be of the form $\begin{pmatrix}G_{h}&A\\ A^{t}&B\end{pmatrix}$, where $G_{h}$ is some positive semidefinite Gram-like matrix of $h$ (cf. Lemma 3.14). Then we provide a way to produce a positive semidefinite matrix out of a partially specified symmetric matrix of the aforementioned structure only by controlling the diagonal elements of $B$ (cf. Lemma 3.17. Exploiting these two results, we further provide some necessary and sufficient conditions in terms of a representation of a SOHS part $h$ of $f$ and monomial decompositions of $f-h$ (cf. Theorem 3.19).
    We provide necessary and sufficient condition for $f=h+\sum_{j}d_{j}\delta_{j}$, with $h$ SOHS and $\delta_{j}$ are not hermitian squares, to have a SOHS extension $\tilde{f}$. If for any decomposition of $\delta_{j}=l_{k}r_{k}$, $l_{k}$ and $r_{k}$ simultaneously do not belong to the set of monomial vectors of a Gram-like matrix $G_{h}$ and if $G_{h}$ is positive definite, we can always extend $f$ to a Gram-like matrix preserving $\tilde{f}$. But if $G_{h}$ is psd but not pd, we need to have additional conditions, namely, the column space of $A$ should be sitting inside the column space of $G_{h}$. We see that the problem boils down to choosing diagonal elements of lower diagonal block of a $2\times 2$ block matrix, so that we can make the block matrix positive semidefinite.
    This, in turn, helps us to find noncommutative polynomials $f$ that do not admit any such extension $\widetilde{f}$ (cf. Example 3.13). We then provide necessary and sufficient conditions for the existence of such $\widetilde{f}$, when $G_{\widetilde{f}}$ takes a special form, namely as in (11) (cf. Theorem 3.22). This characterisation helps us to produce a plenty of SOHS polynomials such that their difference is also a SOHS, which is not true in general (cf. Remark 3.23).
    Apart from these existential criteria of such $\widetilde{f}$ in Theorem 3.19 and Theorem 3.22, we also provide a rigorous construction of $G_{\widetilde{f}}$ (and hence $\widetilde{f}$) for a class of noncommutative polynomials $f=h+\sum_{j=1}^{r}a_{j}\zeta_{j}$, $h$ being SOHS. We do so whenever ${\rm rc}(\eta,i)\neq\zeta_{j}$, $1\leq j\leq r$, or equivalently, whenever the monomials $\zeta_{j}$ are not equal to any right end part of any monomials $\eta$ of $h$ (cf. Theorem 3.24).
  2. (2)
    In subsection 3.3, we deal with Question 1 via graph theory. In fact, we look at positive semidefinite completion problem in the context of completion of a quasi SOHS polynomial (cf. Definition 3.34 for details). We prove the following theorem which describes a pattern made out of the known coefficients of a quasi SOHS polynomial that ensures a SOHS completion of the same (cf. Theorem 3.35). The pattern is given in terms of a graph associated to a partial Gram-like matrix of the given quasi SOHS polynomial. To be specific, we prove the following :
    Let $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ be a quasi SOHS polynomial w.r.t a representation $\mathbf{R}_{f}$, as in (8), along with conditions as in (9). Then any such $f$ has a SOHS completion $\overline{f}^{\;G_{f},\;W_{k}}$ w.r.t the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$ if and only if the specification graph of $G_{f}$ is chordal. Here $W_{k}$ is the column vector $\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}\end{pmatrix}^{t}$, $\zeta_{i}$, $1\leq i\leq k$, being the monomials appearing in $\mathbf{R}_{f}$.
    Here a positive semidefinite Gram-like matrix of any such SOHS completion keeps the positive semidefinite Gram-like matrices of SOHS parts of the given quasi SOHS polynomial as its principal submatrices. We also prove a similar statement for Gram matrix as well, (cf. $(1)$ of Theorem 3.35).
  3. (3)
    In subsection 3.4, we again answer when a quasi SOHS polynomial can be completed to a SOHS but now the pattern of known and unknown entries is guided by a projective algebraic set. Moreover, the defining equations of this algebraic set are homogeneous square-free quadratic (commutative) monomials. The nobleness of this elucidation is that a SOHS completion of some partial symmetric noncommutative polynomial of any even degree is controlled by some homogeneous (commutative) monomials of degree $2$. To be more specific, we prove that $2$-regularity of such a projective algebraic set guarantees the SOHS completion. Furthermore, it is deeply related to Theorem 3.35 as well. Precisely, we obtain the following (cf. Theorem 3.47 for details):
    Let $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ be a quasi SOHS polynomial w.r.t a representation $\mathbf{R}_{f}$, as in (8), along with conditions as in (9). Then any such $f$ has a SOHS completion $\overline{f}^{\;G_{f},\;W_{k}}$ w.r.t the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$ if and only if the subspace arrangement $V_{G_{f}}(\subseteq\mathbb{P}^{k-1})$ of $G_{f}$ is $2$-regular.
    We provide some examples to display simultaneous applications of Theorem 3.35 and Theorem 3.47. In these examples we use Macaulay2 software to calculate regularity via Betti tables and free resolutions.
2. Background material
We introduce our notations and recall necessary definitions and results in this section.
2.1. Noncommutative polynomial and SOHS
Let $X_{1},X_{2},\dots,X_{n}$ be noncommuting variables. Generating all possible words, also referred as monomials, in these letters with finite length by concatenating the letters and adding $1$, the empty word, we get a free monoid. Let $\langle\underline{X}\rangle=\langle X_{1},X_{2},\dots,X_{n}\rangle$ be the free monoid generated by $X_{1},X_{2},\dots,X_{n}$. We take all possible finite linear combinations of words from $\langle\underline{X}\rangle$ with coefficients from $\mathbb{R}$ and we call them real noncommutative polynomials. Denote the set of all real noncommutative polynomials as
$\mathbb{R}\langle\underline{X}\rangle=\mathbb{R}\langle X_{1},X_{2},\dots,X_{n}\rangle=\Big{\{}\sum_{i=1}^{N}a_{i}W_{i}\mid\;a_{i}\in\mathbb{R},N\in\mathbb{N},W_{i}\in\langle\underline{X}\rangle\Big{\}}.$
$\mathbb{R}\langle\underline{X}\rangle$ forms a free algebra. We equip $\mathbb{R}\langle\underline{X}\rangle$ with the * operation
$\mathbb{R}\langle\underline{X}\rangle$ $\longrightarrow\mathbb{R}\langle\underline{X}\rangle$
$f$ $\longmapsto f^{*}.$
Under this map, $\mathbb{R}\cup\{\underline{X}\}$ gets fixed and for given
$W$ $=X_{1}X_{2}\dots X_{n-1}X_{n},$
$W^{*}$ $=X_{n}X_{n-1}\dots X_{2}X_{1}.$
By the degree of a monomial $W$, denoted by ${\rm deg}(W)$, we mean the length of $W$. For example, ${\rm deg}(X_{1}X_{2}^{3}X_{3})={\rm deg}(X_{1}X_{2}X_{2}X_{2}X_{3})=5$. By the degree of a noncommutative polynomial $f$, denoted by ${\rm deg}(f)$, we mean the degree of any one of its highest degree monomials.
We reserve the notation ${\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ to denote the set of all symmetric polynomials of $\mathbb{R}\langle\underline{X}\rangle$, that is,
${\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle=\{f\in\mathbb{R}\langle\underline{X}\rangle\mid\;f=f^{*}\}.$
Let $\langle\underline{X}\rangle_{d}$ be the set of all words in $\langle\underline{X}\rangle$ of at most degree $d$. Let $\mathbb{R}\langle\underline{X}\rangle_{d}$ be the set of all polynomials with monomials coming from $\langle\underline{X}\rangle_{d}$. The cardinality of $\langle\underline{X}\rangle_{d}$ is $s(d,n)=\sum_{i=0}^{d}n^{i}$, where $n$ is the number of noncommuting variables in $\langle\underline{X}\rangle$. Let $W_{s(d,n)}$ be the $s(d,n)\times 1$ column vector of the words in $\langle\underline{X}\rangle_{d}$ taken in the lexicographic order. Given a polynomial ${f}\in\mathbb{R}\langle X\rangle_{d}$, $\mathbf{f}\Big{(}=({f_{w}})_{w\in\mathbb{R}\langle X\rangle_{d}}\Big{)}$ be its coefficient vector. Every polynomial ${f}$ of degree $d$ can be expressed as $f=\sum_{w\in\langle\underline{X}\rangle_{d}}f_{w}w={\mathbf{f}}^{T}W_{s(d,n)}=W_{s(d,n)}^{*}\mathbf{f}$.
Definition 2.1.
A noncommunicative polynomial $f\in\mathbb{R}\langle X\rangle$ is positive if $f(A_{1},\dots,A_{n})$ is positive semidefinite for all tuples $(A_{1},\dots,A_{n})$ of real symmetric matrices of the same order.
Definition 2.2.
For any $p\in\mathbb{R}\langle\underline{X}\rangle$, the noncommutative polynomial $p^{*}p$ is called a Hermitian square. A noncommutative polynomial $f\in\mathbb{R}\langle\underline{X}\rangle$ is said to be a sum of Hermitian squares (or SOHS in short) if it can be written as a sum of some Hermitian squares, i.e. there exists noncommutative polynomials $g_{1},g_{2},\dots,g_{k}\in\mathbb{R}\langle X\rangle$ for some $k\in\mathbb{N}$ such that
$f=g_{1}^{*}g_{1}+g_{2}^{*}g_{2}+\dots+g_{k}^{*}g_{k}.$
Remark 2.3.
A noncommutaive polynomial is positive if and only if it is SOHS. See [16] for Helton’s proof of the result and see [22] for McCullough and Putinar’s alternative proof.
We now recall a notion that formalizes any right end part of a given monomial (cf. [20, Section 3, p. 745] or [9, Definition 2.5, p. 38]).
Definition 2.4.
The right chip function ${\rm rc}:\langle\underline{X}\rangle\times\mathbb{N}\rightarrow\langle\underline{X}\rangle$ is defined by
${\rm rc}(W_{1}W_{2}\cdots W_{n},i)=\left\{\begin{array}[]{ll}W_{n-i+1}W_{n-i+2}\cdots W_{n}&\mbox{if $1\leq i<n$};\\ \\ W_{1}W_{2}\cdots W_{n}&\mbox{if $i\geq n$};\\ \\ 1&\mbox{if $i=0$}.\end{array}\right.$
Here ${\rm deg}(W_{i})=1$ for all $1\leq i\leq n$.
Example 2.5.
Let $W=X_{1}^{2}X_{2}^{2}X_{3}X_{2}^{2}X_{1}$ be a noncommutative word. Then, ${\rm rc}(W,4)=X_{3}X_{2}^{2}X_{1}$, ${\rm rc}(W,8)={\rm rc}(W,9)=X_{1}^{2}X_{2}^{2}X_{3}X_{2}^{2}X_{1}$ and ${\rm rc}(W,0)=1$.
2.2. Some properties of positive definite and semidefinite matrices
We now recall the characterisations of positive definite and positive semidefinite matrices in terms of their principal minors.
Lemma 2.6 (Sylvester’s criterion for positive definite matrices).
A real symmetric matrix is positive definite if all its leading principal minors are positive.
Proof.
See [2, §3.3, 3.15, p. 27]. ∎
Lemma 2.7 (Generalized Sylvester’s criterion for positive semidefinite matrices).
A real symmetric matrix is positive semidefinite if all its principal minors are nonnegative.
Proof.
See [5, §1.2, p. 7]. ∎
The following well-known results are easy consequences of spectral decomposition of a real symmetric matrix and will be used often in the proofs of the results in Section 3. The following results are direct consequences of the spectral theorem for real symmetric matrices.
Lemma 2.8.
Let $G$ be a positive semidefinite matrix with real entries of size $k$. Then the matrix obtained from $G$ by first interchanging its $i$-th and $j$-th row (column respectively) followed by interchanging its $i$-th and $j$-th column (row respectively) is again a positive semidefinite matrix, for any $1\leq i,j\leq k$.
Lemma 2.9.
Let $A$ be a block diagonal matrix with $A_{1},A_{2},\ldots,A_{n}$ being its diagonal blocks. If each $A_{i}$ is a positive definite (respectively positive semidefinite) matrix, then so is $A$.
We recall another well-known property of a positive semidefinite matrix. We present a rudimentary argument, which is available in the literature, for the reader not familiar with this result.
Lemma 2.10.
Let $A$ be a $k\times k$ positive semidefinite matrix with real entries. Let $(i,i)$-th entry of $A$ be zero, for some $1\leq i\leq k$. Then the $i$-th column and the $i$-th row of $A$ is zero.
Proof.
Follows from Lemma 2.7. Indeed, suppose the contrary. Then there will be some $1\leq j\leq k$ such that $(i,j)$-th entry, and hence $(j,i)$-th entry of $A$ is nonzero and both are equal. Then, the determinant of the $2\times 2$ principal submatrix of $A$ formed by $(i,i)$-th entry, $(i,j)$-th entry, $(j,i)$-th entry and $(j,j)$-th entry is negative, which contradicts Lemma 2.7. ∎
We end this subsection by recalling a couple more well-known characterizations of a positive semidefinite matrix in terms of a given positive semidefinite principal minor and its Schur complement.
Lemma 2.11.
Let $M$ be a symmetric matrix with real entries of the following form
$M=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}A\end{matrix}&B\\ \hline\cr B^{t}&\begin{matrix}C\end{matrix}\end{array}\right)\;,$
where $A$ and $C$ are square matrices. Then $M$ is positive semidefinite (respectively positive definite) if and only if $A$ is positive definite and $M/A(:=C-B^{t}A^{-1}B)$ is positive semidefinite (respectively positive definite).
Proof.
See [27, Theorem 1.12, p. 34]. ∎
In fact, this lemma has a generalisation. By $A^{-}$ we denote a generalised inverse of a given square matrix $A$, i.e, $AA^{-}A=A$. Given a matrix $B$, not necessarily square, by $\mathcal{C}(B)$ we denote its column space.
Lemma 2.12.
Let $M$ be a symmetric matrix with real entries of the following form
$M=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}A\end{matrix}&B\\ \hline\cr B^{t}&\begin{matrix}C\end{matrix}\end{array}\right)\;,$
where $A$ and $C$ are square matrices. Then $M$ is positive semidefinite if and only if $A$ is positive semidefinite, $\mathcal{C}(B)\subseteq\mathcal{C}(A)$, and $M/A(:=C-B^{t}A^{-}B)$ is positive semidefinite.
Proof.
See [27, Theorem 1.20, p. 44]. ∎
2.3. Gram Matrix of a symmetric polynomial and a SOHS
Though it is a well known fact that any symmetric noncommutative polynomial has a symmetric matrix associated to it, known as Gram matrix, we still include it in a bit detail as we need this later.
Theorem 2.13.
Any $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ of degree $2d$ in $n$ variables has a representation $\mathbf{R}_{f}$ given as follows:
$\mathbf{R}_{f}\;:\;f=\sum_{i=1}^{s(d,n)}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq s(d,n)}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$ (1)
where
$\begin{split}&s(d,n)=\sum_{k=0}^{d}n^{k}\;,\\ &a_{i},b_{ij}\in\mathbb{R}\;,\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle,\;{\rm deg}(\zeta_{i})\leq d,\;\zeta_{i}\neq\zeta_{j}\text{\;for\;all\;}1\leq i<j\leq s(d,n)\;.\end{split}$ (2)
Hence, there exists a $s(d,n)\times s(d,n)$ symmetric matrix $G_{f}$ such that
$f=W_{s(d,n)}^{\ast}\;G_{f}\;W_{s(d,n)},$ (3)
where $W_{s(d,n)}$ is the column vector of all monomials of degree at most $d$ taken in lexicographic order.
Proof.
Firstly, it is easy to note that any degree $2k$ monomial can be written as a product of two degree $k$ monomials. Now, let $a\eta$ be a term of $f$. As $f$ is chosen from ${\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$, then either $\eta^{\ast}=\eta$ or $a\eta^{\ast}$ must be another term of $f$. Hence, the first part of the assertion follows.
Let $1=\zeta_{1}<\zeta_{2}<\cdots<\zeta_{s(d,n)}=X_{n}^{d}$ be all such monomials in ascending order, where we denote the lexicographic order in the set of all monomials by $<$. Then we have,
$W_{s(d,n)}=\left(\begin{array}[]{@{}c}\begin{matrix}\zeta_{1}&\zeta_{2}&\cdots&\zeta_{s(d,n)}\end{matrix}\end{array}\right)^{t}.$
Choosing
$G_{f}=\begin{pmatrix}a_{1}&b_{12}&b_{13}&\cdots&b_{1s(d,n)}\\ b_{12}&a_{2}&b_{23}&\cdots&b_{2s(d,n)}\\ b_{13}&b_{23}&a_{3}&\cdots&b_{3s(d,n)}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ b_{1s(d,n)}&b_{2s(d,n)}&b_{3s(d,n)}&\cdots&a_{s(d,n)}\end{pmatrix}\;,$ (4)
we have the second part of the assertion. ∎
Definition 2.14.
Let $f$ be any symmetric noncommutative polynomial. The matrix $G_{f}$, as in (4), is called a Gram matrix of $f$. It is also called the Gram matrix associated with the representation $\mathbf{R}_{f}$, as in (1), along with conditions as in (2).
Remark 2.15.
  1. (1)
    It can be noted that given a degree $2d$ symmetric noncommutative polynomial, it doesn’t necessarily have unique representation. For example, consider $f=5X_{1}^{6}+8X_{2}^{4}$. Then $f$ has representations $5(X_{1}^{3})^{\ast}X_{1}^{3}+4(X_{2})^{\ast}X_{2}^{3}+4(X_{2}^{3})^{\ast}X_{2}$ and $5(X_{1}^{3})^{\ast}X_{1}^{3}+4(X_{2}^{2})^{\ast}X_{2}^{2}+4(X_{2}^{2})^{\ast}X_{2}^{2}$.
  2. (2)
    However, the moment we fix a representation of a degree $2d$ symmetric noncommutative polynomial, associated Gram matrix and monomial vector get fixed.
It is easy to note that any SOHS is definitely a symmetric noncommutative polynomial. We want to look at the converse. The following theorem gives a criteria for a symmetric polynomial to be a SOHS in terms of its Gram matrix.
Theorem 2.16.
A symmetric polynomial $f$ of degree $2d$ in noncommuting variables is a SOHS if and only if there exists a positive semidefinite matrix $G_{f}$ such that
$f=W_{s(d,n)}^{\ast}\;G_{f}\;W_{s(d,n)},$ (5)
where $W_{s(d,n)}$ is the column vector consisting of all words in $\langle\underline{X}\rangle$ of degree $\leq d$ taken in lexicographic order.
Proof.
See [11, §2, p. 105-110] or [9, Proposition 1.16, p. 6]. ∎
Remark 2.17.
To express a symmetric polynomial of degree $2d$ in the form as in (3), it is not always necessary to take all monomials of degree at most $d$ in $W_{s(d,n)}$. For example, consider the polynomial $f=1+2X_{1}+X_{1}^{2}+X_{1}X_{2}^{2}+2X_{2}^{2}+X_{2}^{2}X_{1}+X_{2}X_{1}^{2}X_{2}+X_{2}^{4}$. Then for
$W_{5}=\begin{pmatrix}1&X_{1}&X_{1}X_{2}&X_{2}&X_{2}^{2}\end{pmatrix}^{t}\;,$
and
$G=\begin{pmatrix}1&1&0&0&0\\ 1&1&0&0&1\\ 0&0&1&0&0\\ 0&0&0&2&0\\ 0&1&0&0&1\end{pmatrix}\;,$
we have $f=W_{5}^{\ast}\;G\;W_{5}$. But the following Gram matrix of $f$
$G_{f}=\begin{pmatrix}1&1&0&0&0&0&0\\ 1&1&0&0&0&0&1\\ 0&0&0&0&0&0&0\\ 0&0&0&1&0&0&0\\ 0&0&0&0&2&0&0\\ 0&0&0&0&0&0&0\\ 0&1&0&0&0&0&1\end{pmatrix}$
is of much bigger size.
As the choice of the monomial vector $W_{s(d,n)}$ is very strict in Theorem 2.13 and Theorem 2.16, there are several techniques to throw out the unnecessary monomials and making the matrix of smaller size. For example Newton chip method, augmented Newton chip method etc (cf. [9, Section 2.3 & 2.4]) has been used to reduce the size of the column vector $W_{s(d,n)}$, as in (5), and hence the size of matrix $G_{f}$. Moreover, while putting the monomials in $W_{s(d,n)}$, usually lexicographic order is followed. But as long as enough monomials is being collected in the monomial vector to express a given SOHS $f$, Lemma 2.8 says that the order doesn’t matter. We demonstrate this through the following example.
Example 2.18.
Consider $f=X_{1}^{2}+4X_{1}X_{2}+4X_{2}X_{1}+16X_{2}^{2}$. Then, for $(W^{\prime}_{2})^{\ast}=\begin{pmatrix}X_{1}&X_{2}\end{pmatrix}$, $(W^{\prime\prime}_{2})^{\ast}=\begin{pmatrix}X_{2}&X_{1}\end{pmatrix}$, $G^{\prime}=\begin{pmatrix}1&4\\ 4&16\end{pmatrix}$ and $G^{\prime\prime}=\begin{pmatrix}16&4\\ 4&1\end{pmatrix}$, we have $(W^{\prime}_{2})^{\ast}\;G^{\prime}\;W_{2}^{\prime}=f=(W^{\prime\prime}_{2})^{\ast}\;G^{\prime\prime}\;W_{2}^{\prime\prime}$.
It is evident from Remark 2.17 and Example 2.18 that to be able to reduce the size of the monomial vector (and hence that of a Gram matrix) whenever possible and write the monomial entries in that monomial vector in any arbitrary order, while expressing a symmetric or SOHS polynomial in the required form, puts us in a resilient situation. Therefore, we formalize all of this in the upcoming section after putting things in a particular perspective.
3. Extending a noncommutative polynomial to a SOHS
In this section, whenever we consider a noncommutative polynomial $f\in\mathbb{R}\langle\underline{X}\rangle$, where $\langle\underline{X}\rangle=\langle X_{1},X_{2},\dots,X_{n}\rangle$, we mean that all the variables $X_{i}$, $1\leq i\leq n$, are essential to express $f$. Moreover, we want to find a SOHS extension $\widetilde{f}$ of $f$ in $\mathbb{R}\langle\underline{X}\rangle$ only. Specifically, we now prepare the ground for defining Gram-like matrices, which will help us answer the Question 1 that captures the main theme of this article.
3.1. Gram-like matrices for symmetric and SOHS noncommutative polynomials
Recall that Theorem 2.16 provides a way of getting a SOHS polynomial of degree $2d$ in $n$ variables whenever we have a $s(d,n)\times s(d,n)$ positive semidefinite matrix. We observe that a positive semidefinite matrix of arbitrary size also produces a SOHS of correct degree.
Proposition 3.1.
Let $G$ be a $k\times k$ positive semidefinite matrix and $W_{k}$ be a column vector of size $k$ with $k$ many distinct monomial entries (not necessarily in lexicographical order). Then
  1. (1)
    The polynomial $W_{k}^{\ast}\;G\;W_{k}$ is a SOHS polynomial.
  2. (2)
    Moreover, if no row (and therefore column) of $G$ is zero, then degree of the polynomial $W_{k}^{\ast}\;G\;W_{k}$ is twice the degree of any of the highest degree monomials in $W_{k}$.
Proof.
  1. (1)
    Let $d$ be the highest degree among the degrees of the constituent monomials of $W_{k}$. Let $W_{s(d,n)}$ be the monomial vector formed by taking the remaining $s(d,n)-k$ many monomials of degree at most $d$ along with the monomials that are already present in $W_{k}$ and then rearranging them in lexicographic order. This rearrangement is nothing but a permutation between $\{1,2,\cdots,s(d,n)\}$ to itself and we denote it by $\sigma\in S_{s(d,n)}$, the symmetric group on $s(d,n)$ many symbols. Let $G^{\prime}$, a $s(d,n)\times s(d,n)$ matrix, be defined as follows:
    $G^{\prime}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\mbox{\Large 0}\end{array}\right)\;.$
    Let ${G^{\prime}}^{\sigma}$, another $s(d,n)\times s(d,n)$ matrix, be obtained from $G^{\prime}$ by permuting its rows and columns by $\sigma$. Then ${G^{\prime}}^{\sigma}$ is positive semidefinite as $G$ is so and by Lemma 2.9 & Lemma 2.8. Moreover, we have the following equality:
    $W_{k}^{\ast}\;G\;W_{k}=W_{s(d,n)}^{\ast}\;{G^{\prime}}^{\sigma}\;W_{s(d,n)}\;.$
    Therefore, the assertion follows from Theorem 2.16.
  2. (2)
    Let $\zeta_{i_{0}}$ be one of the monomials in $W_{k}$ of maximum degree. Then, by hypothesis, the $i_{0}$-th column and $i_{0}$-th row is nonzero. Therefore, by Lemma 2.10, the $(i_{0},i_{0})$-th entry of $G$ is positive, say $a_{i_{0}}$. Then, $f$ contains the term $a_{i_{0}}\zeta_{i_{0}}^{\ast}\zeta_{i_{0}}$ of degree $2{\rm deg}(\zeta_{i_{0}})$. So, ${\rm deg}(f)\geq 2{\rm deg}(\zeta_{i_{0}})$. The inequality ${\rm deg}(f)\leq 2{\rm deg}(\zeta_{i_{0}})$ is obvious as $W_{k}$ contains no monomial of degree bigger than ${\rm deg}(\zeta_{i_{0}})$. Hence, the assertion follows.
Corollary 3.2.
Let $G$ be any $k\times k$ symmetric matrix and $W_{k}$ be any column vector of size $k$ with $k$ many distinct monomial entries (not necessarily in lexicographical order). Then the polynomial $W_{k}^{\ast}\;G\;W_{k}$ is a symmetric polynomial.
Proof.
Proof goes along the same line of the proof of the first part of the assertion of Proposition 3.1. ∎
Remark 3.3.
  1. (1)
    For a positive definite matrix $G$, every diagonal entry, being a principal minor of $G$, is positive. Therefore, no column or row of $G$ is zero by Lemma 2.10. So, the degree of the SOHS $W_{k}^{\ast}\;G\;W_{k}$ is twice the degree of any of the highest degree monomials in $W_{k}$.
  2. (2)
    Assertion $(2)$ of Proposition 3.1 doesn’t hold for a symmetric matrix in general. Let us take
    $W_{4}:=\begin{pmatrix}1&X_{1}&X_{2}&X_{2}^{2}X_{3}^{2}\end{pmatrix}^{t}\;,$
    and
    $G:=\begin{pmatrix}5&0&0&0\\ 0&1&0&3\\ 0&0&1&0\\ 0&3&0&0\end{pmatrix}\;.$
    Then the polynomial $f=W_{4}^{\ast}\;G\;W_{4}=X_{1}^{2}+X_{2}^{2}+3X_{1}X_{2}^{2}X_{3}^{2}+3X_{3}^{2}X_{2}^{2}X_{1}+5$ is a symmetric polynomial of degree $5$ and the highest degree monomial $X_{2}^{2}X_{3}^{2}$ has degree $4$.
Let $W_{k}$ be a monomial vector with degree of its monomial entries being at most $d$ and $G$ be a $k\times k$ symmetric matrix. Then, as observed in $(2)$ of Remark 3.3, the symmetric polynomial $W_{k}^{\ast}\;G\;W_{k}$ might have degree much lesser than $2d$, unlike the positive semidefinite case. But it is not so under some mild conditions on the symmetric matrix involved.
Proposition 3.4.
Let $G$ be a symmetric matrix with positive diagonal entries and $W_{k}$ be a column vector of distinct monomials from $\mathbb{R}\langle\underline{X}\rangle$. Then the polynomial $f:=W_{k}^{\ast}\;G\;W_{k}$ is a symmetric polynomial of degree $2d$, $d$ being the degree of any of the highest degree monomial entries of $W_{k}$.
Proof.
Firstly, such an $f$ is a symmetric polynomial by Corollary 3.2. Now let $\zeta_{i}$ be a monomial entry of $W_{k}$ having highest degree, say $d$, among all the entries of $W_{k}$ for some $1\leq i\leq k$. Let $a_{i}(>0)$ be the $(i,i)$-th entry of $G$. Then $a_{i}\zeta^{\ast}\zeta$ is a term of $f$ and hence ${\rm deg}(f)\geq 2d$. The other inequality is obvious as no monomial entry of $W_{k}$ has degree more than $d$. ∎
We now record the upshot of the findings of this subsection through a couple of theorems.
Theorem 3.5.
  1. (1)
    Let $G$ be a $k\times k$ positive semidefinite matrix with no row (or column) being zero and $W_{k}$ be a column vector of size $k$ with $k$ many distinct monomial entries of degree at most $d$. Then the polynomial $f=W_{k}^{\ast}\;G\;W_{k}$ is a SOHS polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$ given as follows :
    $\mathbf{R}_{f}\;:\;f=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$ (6)
    where
    $a_{i},b_{ij}\in\mathbb{R},\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle,\;{\rm deg}(\zeta_{i})\leq d\text{\;for\;all\;}1\leq i<j\leq k\;,$ (7a)
    $\zeta_{i}\neq\zeta_{j}\text{\;for\;all\;}1\leq i\neq j\leq k\;,$ (7b)
[…]
    for any $\begin{split}&\text{for\;any\;}K=\{i_{0},i_{1},\ldots,i_{\mid K\mid}\}\subset\{1,2,\ldots,k\}\text{\;with\;}i_{0}<i_{1}<\cdots<i_{\mid K\mid}\\ &\text{the\;matrix\;}\begin{pmatrix}a_{i_{1}}&b_{i_{1}i_{2}}&\cdots&b_{i_{1}i_{\mid K\mid}}\\ b_{i_{1}i_{2}}&a_{i_{2}}&\cdots&b_{i_{2}i_{\mid K\mid}}\\ \vdots&\vdots&\ddots&\vdots\\ b_{i_{1}i_{\mid K\mid}}&b_{i_{2}i_{\mid K\mid}}&\cdots&a_{i_{\mid K\mid}}\end{pmatrix}\text{is\; positive\;semidefinite\;.}\end{split}$ (7d)
  2. (2)
    Conversely, let $f$ be a SOHS polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$, as in (6), satisfying the conditions as given in (7). Then there exists a $k\times k$ positive semidefinite matrix $G$ with no row (or column) being zero and a column vector $W_{k}$ of size $k$ with distinct monomial entries of degree at most $d$ such that $f=W_{k}^{\ast}\;G\;W_{k}$.
Proof.
The condition as in (7c) ensures that no column or row are full zero, that is,
[…]
by Proposition 3.4 and Lemma 2.10. Moreover, the condition as in (7d) ensures that the matrix whose $(i,j)$-th entries are $b_{ij}$ for all $i\neq j$, and $(i,i)$-th entries are $a_{i}$ for all $i$, is positive semidefinite by Lemma 2.7. Now it follows from Proposition 3.1. ∎
Theorem 3.6.
  1. (1)
    Let $G$ be a $k\times k$ symmetric matrix with positive diagonal entries and $W_{k}$ be a column vector of size $k$ with distinct monomial entries of degree at most $d$. Then the polynomial $f=W_{k}^{\ast}\;G\;W_{k}$ is a symmetric polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$ given as follows :
    $\mathbf{R}_{f}\;:\;f=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$ (8)
    where
    $a_{i},b_{ij}\in\mathbb{R},\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle,\;\zeta_{i}\neq\zeta_{j},\;{\rm deg}(\zeta_{i})\leq d\text{\;for\;all\;}1\leq i<j\leq k\;,$ (9a)
    $a_{i}>0\text{\;for\;all\;}1\leq i\leq k\;,$ (9b)
  2. (2)
    Conversely, let $f$ be a symmetric polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$, as in (8), satisfying the conditions as given in (9). Then there exists a $k\times k$ symmetric matrix $G$ with positive diagonal entries and a column vector $W_{k}$ of size $k$ with distinct monomial entries of degree at most $d$ such that $f=W_{k}^{\ast}\;G\;W_{k}$.
Proof.
The condition as in (9b) ensures that no column or row are full zero. The assertion now follows from Proposition 3.4. ∎
Definition 3.7.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a symmetric polynomial of degree $2d$ in $n$ variables and $k$ be any positive integer satisfying $k\leq s(d,n)$. If there exist a $k\times k$ symmetric matrix $G$ with positive diagonal entries such that $f$ can be written in the form $W_{k}^{\ast}\;G\;W_{k}$ for some column vector $W_{k}$ of monomials of degree at most $d$, then $G$ is called a Gram-like matrix of $f$ (or the Gram-like matrix associated with $\mathbf{R}_{f}$, as in (8), along with conditions as in (9)).
Definition 3.8.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a SOHS polynomial of degree $2d$ in $n$ variables and $k$ be any positive integer satisfying $k\leq s(d,n)$. If there exists a $k\times k$ positive semidefinite matrix $G$ such that $f$ can be written in the form $W_{k}^{\ast}\;G\;W_{k}$ for some column vector $W_{k}$ of monomials of degree at most $d$, then $G$ is called a positive semidefinite Gram-like matrix of $f$ (or the positive semidefinite Gram-like matrix associated with $\mathbf{R}_{f}$, as in (6), along with conditions as in (7)).
Remark 3.9.
  1. (1)
    The difference between a Gram matrix and a Gram-like matrix of a polynomial can be seen through the following example. A positive semidefinite Gram-like matrix of the polynomial $f=X_{1}^{2}+4X_{1}X_{2}+4X_{2}X_{1}+16X_{2}^{2}$, as in Example 2.18, is $G^{\prime\prime}=\begin{pmatrix}16&4\\ 4&1\end{pmatrix}$, whereas each of its positive semidefinite Gram matrix is of size $7$.
  2. (2)
    It can be noted from the proof of Proposition 3.1 that a positive semidefinite Gram-like matrix associated to a SOHS polynomial is equally important to its positive semidefinite Gram matrix as it encodes all the information of its Gram matrix. Moreover, it is easier to handle a Gram-like matrix as it usually has much lesser size than corresponding Gram matrix.
  3. (3)
    The motivation for working with a Gram-like matrix of a noncommutative polynomial $f$ instead of its Gram matrix is to avoid unnecessary monomials in the corresponding monomial vector, which is same as ignoring any full zero row (or column) from a Gram matrix. The transition from a Gram matrix to a Gram-like matrix can be interpreted through the lens of graph theory as well. To be specific, the sparsity pattern (cf. [25, §8.1, p. 329-330] for more details) of a Gram-like matrix is always connected, whereas it is not so for a Gram matrix in general.
3.2. Gram-like matrix preserving SOHS extensions
In this subsection, we answer Question 1 completely by exploiting the flexibility obtained through Theorem 3.5 and Theorem 3.6. To do so, let us formally define the meaning of SOHS extension of a noncommutative polynomial.
Definition 3.10.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial of the form $f=\sum_{j=1}^{m}a_{j}\delta_{j}$, where $\delta_{j}$ are monomials in $\mathbb{R}\langle\underline{X}\rangle$, $a_{j}\in\mathbb{R}-\{0\}$ for all $j$. A SOHS polynomial $\widetilde{f}\in\mathbb{R}\langle\underline{X}\rangle$ is called a nontrivial SOHS extension of $f$ if $\widetilde{f}-f=\sum_{i=1}^{n}b_{i}\alpha_{i}$, where $\alpha_{i}$ are monomials in $\mathbb{R}\langle\underline{X}\rangle$, $\alpha_{i}\neq\delta_{j}$, $b_{i}\in\mathbb{R}-\{0\}$ for all $i,j$.
From now on, by a SOHS extension we mean a nontrivial SOHS extension only, as defined above. We now define a special type of SOHS extension, namely, a Gram-like matrix preserving extension.
Definition 3.11.
A SOHS extension $\widetilde{f}\in\mathbb{R}\langle\underline{X}\rangle$ of a noncommutative polynomial $f\in\mathbb{R}\langle\underline{X}\rangle$ is said to be a Gram-like matrix preserving extension of $f$ w.r.t $h$ if
  • there exists a $l\times l$ positive semidefinite Gram-like matrix $G_{\widetilde{f}}$ of $\widetilde{f}$ that keeps a $k\times k$ positive semidefinite Gram-like matrix $G_{h}$ of a SOHS part $h$ of $f$ as a principal submatrix, for some positive integers $l>k$, and
  • there exists a monomial vector $W_{l}$ associated to $G_{\widetilde{f}}$ and a monomial vector $W_{k}$ associated to $G_{h}$ such that $W_{l}$ contains all the entries of $W_{k}$.
Remark 3.12.
From now on, we simply say that $G_{\widetilde{f}}$ is a positive semidefinite Gram-like matrix of a Gram-like matrix preserving extension $\widetilde{f}$ of $f$ to mean that such $G_{\widetilde{f}}$ satisfies both the conditions as in Definition 3.11.
Example 3.13.
  1. (1)
[…]
  2. (2)
    By definition, it is obvious that any Gram-like matrix preserving extension w.r.t some SOHS part of a given noncommutative polynomial is a SOHS extension. However, the converse is not true. Consider the polynomial $f=X^{2}+X+5$ with $h=X^{2}+5$. Then $\widetilde{f}=X^{2}+X+5+X^{4}=f+X^{4}$ is a SOHS extension of $f$. But, as explained above, $f$ doesn’t admit a Gram-like matrix preserving extension w.r.t $h$.
The following lemma talks about the general form of a positive semidefinite Gram-matrix of a Gram-like matrix preserving extension.
Lemma 3.14.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial and $h$ a SOHS part of $f$. Let $\widetilde{f}$ be a Gram-like matrix preserving extension of $f$ w.r.t $h$ and $G_{h}$ and $G_{\widetilde{f}}$ be as in Definition 3.11. Then, up to permutation, $G_{\widetilde{f}}$ must be of the form
$G_{\widetilde{f}}=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&A\\ \hline\cr A^{t}&\begin{matrix}B\end{matrix}\end{array}\right)\;,$
where $G_{\widetilde{f}}/A(:=A^{t}B^{-}A)$ is positive semidefinite and $\mathcal{C}(A)\subseteq\mathcal{C}(G_{h})$.
Proof.
Follows from Definition 3.11, Lemma 2.8 and Lemma 2.12. ∎
We now look for some necessary conditions and sufficient conditions for the existence of a Gram-like matrix preserving extension of a given noncommutative polynomial $f$, in terms of the monomials of $f$. We set the stage for that.
The following lemma talks about how a given monomial can be decomposed through two smaller sized monomials.
Lemma 3.15.
Any monomial $W\in\langle\underline{X}\rangle$ can be decomposed as a product of two monomials in ${\rm deg}(W)+1$ many ways.
Proof.
Any monomial $W\in\langle\underline{X}\rangle$ can be decomposed as following:
$W=\big{(}{\rm rc}(W^{*},{\rm deg}(W)-n)\big{)}^{*}{\rm rc}(W,n),$ (10)
for any integer $n$ satisfying $0\leq n\leq{\rm deg}(W)$ and therefore there are ${\rm deg}(W)+1$ many such ways. ∎
Definition 3.16.
Given a monomial $W\in\langle\underline{X}\rangle$, any decomposition of $W$, as in (10), is called a rc deomposition of $W$ (Right Chip decomposition of $W$).
The following lemma provides a way to produce a positive semidefinite matrix out of a partially specified symmetric matrix of a particular type. This is very crucial in proving the existence of a Gram-like matrix preserving extension.
Lemma 3.17.
Let
$G=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}A\end{matrix}&B\\ \hline\cr B^{t}&\begin{matrix}C\end{matrix}\end{array}\right)\;$
be a partially specified symmetric matrix, where $A$ and $C$ are square matrices of size $m$ and $n$ respectively and $A$ is a positive semidefinite matrix. Let the diagonal elements of $C$ be the only unspecified entries of $G$. Then if the column space of $B$ sits inside the column space of $A$, that is, $\mathcal{C}(B)\subseteq\mathcal{C}(A)$, then it is possible to fill the diagonal elements of $C$ with some positive real numbers so that the matrix $G$ becomes positive semidefinite.
Proof.
By Lemma 2.12, it is enough to choose diagonals $c_{ii}$ of $C$ such that
  • $c_{ii}>0$ for all $1\leq i\leq n$,
  • The matrix $G/A:=C-B^{t}A^{-}B$ is positive semidefinite.
We choose $c_{ii}>0$ satisfying $c_{ii}\geq\sum_{j\neq i}|(C-B^{t}A^{-}B)_{ij}|$, for all $i$. Clearly, such choices make the matrix $G/A:=C-B^{t}A^{-}B$ diagonally dominant, and hence positive semidefinite. Hence the assertion follows. ∎
Remark 3.18.
In Lemma 3.17, if the matrix $A$ is positive definite, then condition $\mathcal{C}(B)\subseteq\mathcal{C}(B)$ is automatically satisfied.
We are now ready to provide some necessary and sufficient conditions for the existence of a Gram-like matrix preserving extension.
Theorem 3.19.
Let $f=h+\sum_{j=1}^{m}d_{j}\delta_{j}\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial with a given SOHS part $h$ and the monomials $\delta_{j}$s are not Hermitian squares.
  1. (1)
    Assume that there exists a Gram-like matrix preserving extension $\widetilde{f}$ of $f$ w.r.t $h$. Then, there exists a representation
    $\mathbf{R}_{h}\;:\;h=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})\;,$
    and, for all $1\leq j\leq m$, there exist rc decompositions of $\delta_{j}=\beta_{n_{j}}^{\ast}\gamma_{n_{j}}$, for some $0\leq n_{j}\leq{\rm deg}(\delta_{j})$, such that no $\beta_{n_{j}}$ and $\gamma_{n_{j}}$ simultaneously lie in $\{\zeta_{1},\ldots,\zeta_{k}\}$.
  2. (2)
    Conversely, assume that there exists a representation
    $\mathbf{R}_{h}\;:\;h=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})\;,$
    and, for all $1\leq j\leq m$, there exist rc decompositions of $\delta_{j}=\beta_{n_{j}}^{\ast}\gamma_{n_{j}}$, for some $0\leq n_{j}\leq{\rm deg}(\delta_{j})$, such that no $\beta_{n_{j}}$ and $\gamma_{n_{j}}$ simultaneously lie in $\{\zeta_{1},\ldots,\zeta_{k}\}$. Then, there exists a partially specified symmetric matrix $G$ of the form
    $G=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&A\\ \hline\cr A^{t}&\begin{matrix}B\end{matrix}\end{array}\right)\;,$
    where $G_{h}$ is the Gram-like matrix of $h$ w.r.t the representation $\mathbf{R}_{h}$. Moreover, if $\mathcal{C}(A)\subseteq\mathcal{C}(G_{h})$, then there exist infinitely many Gram-like matrix preserving extensions of $f$.
Proof.
  1. (1)
    As $\widetilde{f}$ is a Gram-like matrix preserving extension w.r.t $h$, by Lemma 3.14 it must be of the form
    $G_{\widetilde{f}}=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&A\\ \hline\cr A^{t}&\begin{matrix}B\end{matrix}\end{array}\right)\;$
    up to a permutation, where $G_{h}$ is a positive semidefinite Gram-like matrix of $h$. Let $G_{h}$ be the Gram-like matrix associated with
    $\mathbf{R}_{h}\;:\;h=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})\;.$
    1. Case 1
      All the coefficients $d_{j}$ of the monomials $\delta_{j}$ of $f-h$ are distributed in $B$ :
      In this case, there exist some $0\leq n_{j}\leq{\rm deg}(\delta_{j})$, such that the decompositions
      $\delta_{j}=({\rm rc}(\delta_{j}^{\ast},{\rm deg}(\delta_{j})-n_{j}))^{\ast}{\rm rc}(\delta_{j},n_{j})$
      satisfies
      ${\rm rc}(\delta_{j}^{\ast},{\rm deg}(\delta_{j})-n_{j})\neq\zeta_{i}\neq{\rm rc}(\delta_{j},n_{j}),$
      for all $1\leq i\leq k$ and $1\leq j\leq m$.
    2. Case 2
      Not all the coefficients $d_{j}$ of the monomials $\delta_{j}$ of $f-h$ are distributed in $B$ :
      In this case, some $d_{j}$ must be distributed in $A$ or $A^{t}$ or both. Therefore,
      1. (i)
        either there exist decompositions $\delta_{j}=({\rm rc}(\delta_{j}^{\ast},{\rm deg}(\delta_{j})-n_{j}))^{\ast}{\rm rc}(\delta_{j},n_{j})$, for some $0\leq n_{j}\leq{\rm deg}(\delta_{j})$, such that
        $\beta_{n_{j}}(:={\rm rc}(\delta_{j}^{\ast},{\rm deg}(\delta_{j})-n_{j}))=\zeta_{i}$
        for some $i\in K_{1}\subseteq\{1,\ldots,k\}$ and $j\in M_{1}\subseteq\{1,\ldots,m\}$. Then, we must have
[…]
        as otherwise these $d_{j}$s would have been distributed inside $G_{h}$, which is not possible as $\widetilde{f}$ is a Gram-like matrix preserving extension of $f$ w.r.t $h$.
      2. (ii)
        Exactly similar to (i).
  2. (2)
    Reversing the implications of the proof the first part of the assertion, it is easy to see that if no $\beta_{n_{j}}$ and $\gamma_{n_{j}}$ simultaneously lie in $\{\zeta_{1},\ldots,\zeta_{k}\}$, then there exists a partially specified symmetric matrix $G$ of the form
    $G=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&A\\ \hline\cr A^{t}&\begin{matrix}B\end{matrix}\end{array}\right)\;,$
    where $G_{h}$ is the Gram-like matrix of $h$ w.r.t the representation $\mathbf{R}_{h}$, such that the coefficient $d_{j}$s of $\delta_{j}$s and $\delta_{j}^{\ast}$s are distributed in $G$ outside the block $G_{h}$. Here we are taking $W_{l}$ as a associated monomial vector to $G$ obtained by taking $\zeta_{i}$s as the first $k$ constituent monomials and taking $\beta_{n_{j}}$s and $\gamma_{n_{j}}$s (without any repetition) as the remaining monomials. Also $d_{j}$s can’t be distributed in the diagonals of $B$, as $\delta_{j}$s are not Hermitian squares.
    Therefore we now have a partially specified symmetric matrix $G=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&A\\ \hline\cr A^{t}&\begin{matrix}B\end{matrix}\end{array}\right)$ such that $G_{h}$ and $B$ are square matrices and the only unspecified entries of $G$ are the diagonal entries of $B$. Now, if $\mathcal{C}(A)\subseteq\mathcal{C}(G_{h})$, then by Lemma 3.17, we can choose diagonal entries of $B$ in such a way that makes $G$ a positive semidefinite matrix. Therefore, the SOHS $\widetilde{f}:=W_{l}^{\ast}\;G\;W_{l}$ is the required Gram-like preserving extension of $f$.
    Moreover, it is clear from Lemma 3.17 that there are infinitely many choices for the diagonal entries of $B$ that makes $G$ positive semidefinite. Therefore, we obtain infinitely many such Gram-like preserving extension $\widetilde{f}$ of $f$.
Remark 3.20.
  1. (1)
    In Theorem 3.19, the hypothesis that the monomials $\delta_{j}$s of $f-h$ are non Hermitian is used to prove the converse part only.
  2. (2)
    Whenever the matrix $G_{h}$ of $h$, as in Theorem 3.19, is positive definite, then the necessary condition itself becomes sufficient as the condition $\mathcal{C}(A)\subseteq\mathcal{C}(G_{h})$ is satisfied by default.
  3. (3)
[…]
    Assume that there exists a Gram-like matrix preserving extension $\widetilde{f}$ of $f$ w.r.t $h$. Because of the possible rc decompositions of $X_{1}$ and $X_{2}$, as mentioned earlier, $\widetilde{f}$ must have a constant term, say $d$. Then for any such $\widetilde{f}$, there should be some positive semidefinite Gram-like matrix $G_{\widetilde{f}}$ such that it contains either $\begin{pmatrix}1&1&1\\ 1&1&-1\\ 1&-1&d\end{pmatrix}$ or $\begin{pmatrix}1&1&-1\\ 1&1&1\\ -1&1&d\end{pmatrix}$, corresponding to monomial vector $\begin{pmatrix}X_{1}&X_{2}&1\end{pmatrix}^{t}$ or $\begin{pmatrix}X_{2}&X_{1}&1\end{pmatrix}^{t}$, as a principal minor. But any of these principal minors are having determinant $-4(<0)$. Therefore, by Lemma 2.7, such $G_{\widetilde{f}}$, and hence such $\widetilde{f}$, can’t exist.
Example 3.21.
[…]
Now we want to look that under what circumstances, a Gram-like matrix preserving extension $\widetilde{f}$ of $f=h+\sum_{j=1}^{m}d_{j}\delta_{j}$ admits a Gram-like matrix $G_{\widetilde{f}}$ of the form $G_{\widetilde{f}}=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}G_{\widetilde{f}-h}\end{matrix}\end{array}\right)$. Here we omit the additional assumption that the monomials $\delta_{j}$s are not Hermitian squares.
Theorem 3.22.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial. Let $h$ be a SOHS part of $f$. Let $\widetilde{f}$ be a SOHS extension of $f$. Then the following are equivalent :
  1. (1)
    $\widetilde{f}$ is a Gram-like matrix preserving extension of $f$ w.r.t $h$ such that there exists a positive semidefinite Gram-like matrix $G_{\widetilde{f}}$ of $\widetilde{f}$ of the form
    $G_{\widetilde{f}}=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}G_{\widetilde{f}-h}\end{matrix}\end{array}\right)\;,$ (11)
    where $G_{h}$ and $G_{\widetilde{f}-h}$ are $k\times k$ and $l\times l$ positive semidefinite Gram-like matrices of $h$ and $\widetilde{f}-h$ respectively.
  2. (2)
    There exist a representation $\mathbf{R}_{h}$ of $h$ and and a representation $\mathbf{R}_{\widetilde{f}-h}$ of $\widetilde{f}-h$, given as
    $\begin{split}\mathbf{R}_{h}\;:\;h=&\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})\;,\\ \mathbf{R}_{\widetilde{f}-h}\;:\;\widetilde{f}-h=&\sum_{i=1}^{l}c_{i}\eta_{i}^{\ast}\eta_{i}+\sum_{1\leq i<j\leq l}d_{ij}(\eta_{i}^{\ast}\eta_{j}+\eta_{j}^{\ast}\eta_{i}),\end{split}$
    both satisfying conditions as in (7), such that
    $\zeta_{i}\neq\eta_{j}\;\text{for\;all\;}1\leq i\leq k,1\leq j\leq l.$
Proof.
$(1)\Rightarrow(2)$ : By Definition 3.11 and Remark 3.12, there exists a monomial vector
$\widetilde{W_{k+l}}:=\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}&\eta_{1}&\cdots&\eta_{l}\end{pmatrix}\;^{t}$
associated to $G_{\widetilde{f}}$ such that $h=W_{k}^{\ast}\;G_{h}\;W_{k}$. Here $W_{k}=\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}\end{pmatrix}\;^{t}$. Then,
$\begin{split}\widetilde{f}-h&=\widetilde{W_{k+l}}^{\ast}\;\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}G_{\widetilde{f}-h}\end{matrix}\end{array}\right)\;\widetilde{W_{k+l}}-W_{k}^{\ast}\;G_{h}\;W_{k}\\ &=W_{l}^{\ast}\;G_{\widetilde{f}-h}\;W_{l}\;,\end{split}$
where $W_{l}=\begin{pmatrix}\eta_{1}&\cdots&\eta_{l}\end{pmatrix}\;^{t}$. Now, let $\mathbf{R}_{h}$ and $\mathbf{R}_{\widetilde{f}-h}$ be the associated representations of $G_{h}$ and $G_{\widetilde{f}-h}$ respectively. Then, $\mathbf{R}_{h}$ and $\mathbf{R}_{\widetilde{f}-h}$ must satisfy the conditions as in (7). Moreover, $\zeta_{i}\neq\eta_{j}$ for all $1\leq i\leq k$, $1\leq j\leq l$, as $\widetilde{W_{k+l}}$ is a monomial vector.
$(2)\Rightarrow(1)$ : Let $G_{h}$ and $G_{\widetilde{f}-h}$ be positive semidefinite Gram-like matrices of $h$ and $\widetilde{f}-h$ associated with representations $\mathbf{R}_{h}$ and $\mathbf{R}_{\widetilde{f}-h}$ respectively. As, $\zeta_{i}\neq\eta_{j}$ for all $1\leq i\leq k$, $1\leq j\leq l$, we can form the the monomial vector
$\widetilde{W_{k+l}}:=\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}&\eta_{1}&\cdots&\eta_{l}\end{pmatrix}\;^{t}.$
Then, denoting $\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}\end{pmatrix}\;^{t}$ and $\begin{pmatrix}\eta_{1}&\cdots&\eta_{l}\end{pmatrix}\;^{t}$ by $W_{k}$ and $W_{l}$ respectively, we have:
$\begin{split}\widetilde{f}&=h+(\widetilde{f}-h)\\ &=W_{k}^{\ast}\;G_{h}\;W_{k}+W_{l}^{\ast}\;G_{\widetilde{f}-h}\;W_{l}\\ &=\widetilde{W_{k+l}}^{\ast}\;\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}G_{\widetilde{f}-h}\end{matrix}\end{array}\right)\;\widetilde{W_{k+l}}\;.\end{split}$
Remark 3.23.
Although the sum $h_{1}+h_{2}$ of two SOHS polynomials $h_{1}$ and $h_{2}$ is always SOHS, the difference $h_{1}-h_{2}$ does not enjoy the same. An example would be $h_{1}=1+X_{1}^{2}+X_{1}X_{2}+X_{2}X_{1}+X_{2}^{2}$ and $h_{2}=1+X_{1}^{2}$, as $h_{1}-h_{2}=X_{1}X_{2}+X_{2}X_{1}+X_{2}^{2}$ is not a SOHS. Comparing this with Example 3.21, we conclude that an important application of Theorem 3.22 is to produce nontrivial SOHS polynomials such that their difference is also a SOHS.
Theorem 3.22 talks about an existential criterion for a Gram-like matrix $G_{\widetilde{f}}$ of the extension $\widetilde{f}$ of $f$ of the form as in (11). We now present one such method of construction of $G_{\widetilde{f}}$ for a class of noncommutative polynomials.
Theorem 3.24.
Let $f=h+\sum_{j=1}^{r}a_{j}\zeta_{j}\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial, without any constant term, where $h$ is a SOHS. Assume that ${\rm rc}(\eta,i)\neq\zeta_{j}$, $1\leq j\leq r$, for any $i\geq 0$ and for any monomial $\eta$ of $h$. Then
  1. (1)
    It is always possible to obtain a Gram-like matrix preserving extension $\widetilde{f}$ of $f$ w.r.t $h$ by adding at most $12r+$ 1 terms with $f$.
  2. (2)
    In fact, infinitely many such $\widetilde{f}$ can be obtained.
Proof.
Let ${\rm deg}(f)=d$. By hypothesis, $f$ can be written as follows:
$f=h+\sum_{j=1}^{r}a_{j}\zeta_{j}=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+\sum_{j=1}^{r}a_{j}\zeta_{j},$
where $a\in\mathbb{R}$, $g_{i}$s and $\zeta$ lie in $\mathbb{R}\langle\underline{X}\rangle$, ${\rm deg}(g_{i})\leq\tfrac{d}{2}$ if $d$ is even or ${\rm deg}(g_{i})\leq\tfrac{d-1}{2}$ if $d$ is odd, for all $1\leq i\leq m$ and $1\leq{\rm deg}(\zeta_{j})\leq d$ for all $1\leq j\leq r$. Let $G_{h}$, say of order $k$, be a positive semidefinite Gram-like matrix of $h$ such that there exist a column vector $W_{k}$ satisfying
  1. (a)
    $1$ does not appear as an entry of $W_{k}$,
  2. (b)
    $\zeta_{1},\ldots,\zeta_{r}$ do not appear as entries of $W_{k}$, and
  3. (c)
    $h=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}=W_{k}^{\ast}\;G_{h}\;W_{k}.$ (12)
It can be noted that existence of such $W_{k}$ satisfying
  • (a) is guaranteed by the assumption that $f$, hence $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$, doesn’t contain any constant term,
  • (b) is guaranteed by the hypothesis ${\rm rc}(\eta,i)\neq\zeta_{j}$, $1\leq j\leq r$, for any $i\geq 0$ and for any monomial $\eta$ of the SOHS part of $f$ (Indeed, let $\zeta_{j}$ appears as an entry of $W_{k}$ for some $j$. Then $\zeta_{j}^{\ast}\zeta_{j}$ appears in the expansion of $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$ as its coefficient in nonzero by (7c) of Theorem 3.5. But then ${\rm rc}(\zeta_{j}^{\ast}\zeta_{j},{\rm deg}(\zeta_{j}))=\zeta_{j}$.),
  • (c) is guaranteed by Theorem 3.5.
Now we consider the matrix $\widetilde{G}$ given as follows:
$\widetilde{G}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}\sum_{j=1}^{r}d_{i}^{2}+1&d_{1}&d_{2}&\cdots&d_{r}\\ d_{1}&1&0&\cdots&0\\ d_{2}&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ d_{r}&0&0&\cdots&1\end{matrix}\end{array}\right)\;,$ (13)
where
$d_{i}=\left\{\begin{array}[]{ll}a_{i}&\mbox{if $\zeta_{i}^{\ast}\neq\zeta_{i}$}\;;\\ \\ \tfrac{a_{i}}{2}&\mbox{if $\zeta_{i}^{\ast}=\zeta_{i}$}\;.\end{array}\right.$
The lower diagonal block of the matrix $\widetilde{G}$, as in (13), is positive semidefinite by Lemma 2.11. Moreover, as the matrix $G_{h}$, as in (12), is positive semidefinite, so is the matrix $\widetilde{G}$ by Lemma 2.9.
We now define $\widetilde{W_{k+r+1}}$ as follows.
$\widetilde{W_{k+r+1}}:=\left(\begin{array}[]{@{}c}\begin{matrix}W_{k}^{t}&1&\zeta_{1}&\cdots&\zeta_{r}\end{matrix}\end{array}\right)^{t}\;.$ (14)
We define $\widetilde{f}$ as follows :
$\widetilde{f}:=\widetilde{W_{k+r+1}}^{\ast}\;\widetilde{G}\;\widetilde{W_{k+r+1}},$ (15)
where $\widetilde{G}$ and $\widetilde{W}_{k+r+1}$ are as in (13) and (14) respectively. Then we have
$\begin{split}\widetilde{f}&=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+\sum_{j=1}^{r}d_{j}\zeta_{j}+\sum_{j=1}^{r}d_{j}\zeta_{j}^{\ast}+\sum_{j=1}^{r}\zeta_{j}^{\ast}\zeta_{j}+\sum_{j=1}^{r}d_{i}^{2}+1\\ &=f+\sum_{\zeta_{j}^{\ast}\neq\zeta_{j}}a_{j}\zeta_{j}^{\ast}+\sum_{j=1}^{r}\zeta_{j}^{\ast}\zeta_{j}+\sum_{j=1}^{r}d_{i}^{2}+1.\end{split}$ (16)
It can be noted that $\widetilde{f}$ is obtained by adding at most $12r+$ 1 many terms with $f$. Therefore, first part of the assertion now follows from Theorem 3.5, (15) and (16).
It can be observed that any $\widetilde{G}$ obtained by replacing the $(k+1,k+1)$-th entry of the matrix by any $A\geq\sum_{j=1}^{r}d_{i}^{2}$ will be positive semidefinite by Lemma 2.11 and hence will produce another $\widetilde{f}$ satisfying the first part of the assertion. Hence, we have infinitely many choices for such $\widetilde{f}$. ∎
3.3. Positive semidefinite completion problem and SOHS via graphs
Positive semidefinite completion problem is one of the most important problems not only in the realm of mathematics but also in physics. Let $A$ be a $n\times n$ partially specified positive semidefinite matrix, i.e, a matrix with some unknown entries and some natural conditions on the known entries. Positive semidefinite completion problem deals with finding the pattern of known entries (and hence of the unknown entries as well) for which $A$ can be completed to a positive semidefinite matrix. Apart from having huge applications, this problem is very intriguing on its own. For example, a solution of this problem has been given via graph theory. It is known that such a solution exists when the pattern of known entries is given by a chordal graph.
In this subsection, we answer Question 1 via positive semidefinite completion problem. We recall some basic definitions and results related to this and study the same problem from the context of extending a partial SOHS polynomial to a SOHS polynomial.
Definition 3.25.
A partially specified $n\times n$ matrix is a partial symmetric matrix if
  • all the diagonal entries are specified, and
  • if the $(i,j)$-th entry is specified, then so is the $(j,i)$-th entry and they are equal.
Definition 3.26.
A partial symmetric matrix is said to be a partial positive definite matrix (respectively partial positive semidefinite matrix) if any of its completely specified principal minor is positive (respectively nonnegative).
Definition 3.27.
The specification graph $\mathcal{G}_{A}$ of a $n\times n$ partial symmetric matrix $A$ is a graph with vertex set $V=\{1,2,\ldots,n\}$ and for $i,j\in V$ with $i\neq j$, the vertices $i$ and $j$ are adjacent whenever $(i,j)$-th entry of $A$ is specified.
Definition 3.28.
A graph $\mathcal{G}$ with vertex set $\{1,2,\ldots,n\}$ is positive definite completable (respectively positive semidefinite completable) if any $n\times n$ partial positive definite (respectively semidefinite) matrix with specification graph $\mathcal{G}$ is completable to a positive definite (respectively semidefinite) matrix.
Definition 3.29.
A graph is said to be chordal if it doesn’t contain the cycle $\mathcal{C}_{k}$ on $k$ vertices as an induced subgraph for any $k\geq 4$.
Example 3.30.
  1. (1)
    The complete graph $\mathcal{K}_{n}$ on $n$ vertices, trees are some examples of chordal graphs.
  2. (2)
    The graph $\mathcal{C}_{4}$, the cycle on $4$ vertices, is not a chordal graph.
Theorem 3.31.
A graph $\mathcal{G}$ is positive definite (semidefinite) completable if and only if $\mathcal{G}$ is chordal.
Proof.
See [15, Theorem 7, p. 120] and [3, Lemma 12.7, p. 161 and Theorem 12.11, p. 163]. ∎
Now we describe what one means by incompleteness of a symmetric noncommutative polynomial w.r.t its partially specified Gram-like matrix and its completion.
Definition 3.32.
A symmetric noncommutative polynomial $f$ of degree $2d$ is said to be incomplete w.r.t its representation $\mathbf{R}_{f}$, as in (8), along with conditions as in (9), if
  • all $a_{i}$s are specified (positive real numbers by conditions as in (9b)), and
  • some $b_{ij}$s are unspecified.
Definition 3.33.
Let $W_{k}$ be a column vector of size $k$ of all monomials from $\mathbb{R}\langle\underline{X}\rangle$ of degree at most $d$ taken in any arbitrary order and $G$ a $k\times k$ partial positive semidefinite matrix with positive diagonal entries. Then for any positive semidefinite completion $\overline{G}$ of $G$, the polynomial $W_{k}^{\ast}\;\overline{G}\;W_{k}$ is called a SOHS completion of the polynomial $f:=W_{s(d,n)}^{\ast}\;G\;W_{s(d,n)}$ w.r.t partial positive semidefinite Gram-like matrix $G$, and is denoted by $\overline{f}^{\;G,\;W_{k}}$.
We now want to look at incomplete polynomials that are not SOHS, but potentially very close to SOHS polynomials. That is, whenever we collect appropriate monomials (with known coefficients) of such an incomplete polynomial $f$ which have the potential to form a SOHS, it does so.
Definition 3.34.
Let $f$ be an incomplete polynomial w.r.t a given representation $\mathbf{R}_{f}$, as in (8), along with conditions as in (9). Then $f$ is said to be quasi SOHS w.r.t $\mathbf{R}_{f}$ if for any $K\subset\{1,2,\ldots,k\}$ such that $b_{ij}$s are known for all $i,j\in K$,
$\sum_{i\in K}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{i,j\in K,i<j}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})$
is a SOHS.
We are now in a stage to reformulate Theorem 3.31 in the context of the dictum of this paper. To be specific, we provide a criteria regarding when a quasi SOHS polynomial qualifies to a SOHS polynomial. We have the following theorem in that regard.
Theorem 3.35.
Let $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ be a quasi SOHS polynomial w.r.t a representation $\mathbf{R}_{f}$, as in (8), along with conditions as in (9). Then any such $f$ has a SOHS completion $\overline{f}^{\;G_{f},\;W_{k}}$ w.r.t the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$ if and only if the specification graph of $G_{f}$ is chordal. Here $W_{k}$ is the column vector $\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}\end{pmatrix}^{t}$, $\zeta_{i}$, $1\leq i\leq k$, being the monomials appearing in $\mathbf{R}_{f}$.
Proof.
Let $f$ be a quasi SOHS polynomial of degree $2d$ w.r.t a representation
$\mathbf{R}_{f}\;\;:\;\;f=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$
where $a_{i},b_{ij}\in\mathbb{R}$, $\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle$ are all monomials of degree at most $d$. Therefore, for any $K\subset\{1,2,\ldots,k\}$ such that $b_{ij}$s are known for all $i,j\in K$,
$\sum_{i\in K}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{i,j\in K,i<j}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})$
is a SOHS. Hence, for any such set $K=\{i_{0},i_{1},\ldots,i_{\mid K\mid}\}$ with $i_{0}<i_{1}<\cdots<i_{\mid K\mid}$, the matrix
$\begin{pmatrix}a_{i_{1}}&b_{i_{1}i_{2}}&\cdots&b_{i_{1}i_{\mid K\mid}}\\ b_{i_{1}i_{2}}&a_{i_{2}}&\cdots&b_{i_{2}i_{\mid K\mid}}\\ \vdots&\vdots&\ddots&\vdots\\ b_{i_{1}i_{\mid K\mid}}&b_{i_{2}i_{\mid K\mid}}&\cdots&a_{i_{\mid K\mid}}\end{pmatrix}$
is positive semidefinite by Theorem 3.5. Therefore any specified principal minor of the Gram-like matrix $G_{f}$, associated with $\mathbf{R}_{f}$ is positive semidefinite, that is to say $G_{f}$ is a partial positive semidefinite matrix. Therefore, it makes sense to talk about SOHS completion of $f$ w.r.t $G_{f}$ following Definition 3.33. Now, the assertion follows from Theorem 3.31. ∎
Remark 3.36.
  1. (1)
    It is easy to observe that Theorem 3.35 holds for Gram matrix as well, by taking $k=s(d,n)$. To do so, one only has to analogously define Definition 3.32, Definition 3.33 and Definition 3.34 w.r.t a representation as $\mathbf{R}_{f}$, as in (1), along with conditions in (2).
  2. (2)
    If we take the specification graph of $G_{f}$ as the complete graphs $\mathcal{K}_{s(d,n)}$ and $\mathcal{K}_{k}$, then we get back Theorem 2.16 and Theorem 3.5 respectively. That is to say, Theorem 3.35 generalises Theorem 2.16 and Theorem 3.5.
  3. (3)
    Let $\mathcal{G}$ be a nonchordal graph. Then also there might exist a quasi SOHS polynomial $f$ with $\mathcal{G}$ as the specification graph of its Gram matrix $G_{f}$ which has a SOHS completion. For example, let $\zeta_{1}=1$, $\zeta_{2}=X_{1}X_{2}$, $\zeta_{3}=X_{2}X_{1}$ and $\zeta_{4}=X_{2}^{2}$ and $W_{4}=\begin{pmatrix}1&X_{1}X_{2}&X_{2}X_{1}&X_{2}^{2}\end{pmatrix}^{t}$. Consider the following polynomial $f$ given as
    $\begin{split}f=&5\zeta_{1}^{\ast}\zeta_{1}+5\zeta_{2}^{\ast}\zeta_{2}+5\zeta_{3}^{\ast}\zeta_{3}++5\zeta_{4}^{\ast}\zeta_{4}+5(\zeta_{1}^{\ast}\zeta_{2}+\zeta_{2}^{\ast}\zeta_{1})+\ast(\zeta_{1}^{\ast}\zeta_{3}+\zeta_{3}^{\ast}\zeta_{1})\\ &+5(\zeta_{1}^{\ast}\zeta_{4}+\zeta_{4}^{\ast}\zeta_{1})+5(\zeta_{2}^{\ast}\zeta_{3}+\zeta_{3}^{\ast}\zeta_{2})+\ast(\zeta_{2}^{\ast}\zeta_{4}+\zeta_{4}^{\ast}\zeta_{2})+5(\zeta_{3}^{\ast}\zeta_{4}+\zeta_{4}^{\ast}\zeta_{3})\\ =5&+(5+\ast)X_{1}X_{2}+(5+\ast)X_{2}X_{1}+10X_{2}^{2}+5X_{1}X_{2}^{2}X_{1}+5X_{2}X_{1}^{2}X_{2}+5X_{1}X_{2}X_{1}X_{2}\\ &+5X_{2}X_{1}X_{2}X_{1}+5X_{1}X_{2}^{3}+5X_{2}^{3}X_{1}+\ast X_{2}X_{1}X_{2}^{2}+\ast X_{2}^{2}X_{1}X_{2}+5X_{2}^{4}.\end{split}$ (17)
    It is easy to check that $f$ is a quasi SOHS polynomial of degree $4$ w.r.t the given representation and the matrix $G_{f}$ given by
    $G_{f}=\begin{pmatrix}5&5&\ast&5\\ 5&5&5&\ast\\ \ast&5&5&5\\ 5&\ast&5&5\\ \end{pmatrix}$
    is a partial positive semidefinite Gram-like matrix of $f$. The matrix $G_{f}$ is also a partial positive semidefinite matrix as $f$ is quasi SOHS and by Theorem 3.5. It can also be noted that the specification graph of $G_{f}$ is $\mathcal{C}_{4}$, which is not chordal. Still there exists a SOHS completion $\overline{f}^{\;G_{f},\;W_{4}}$, which is obtained by taking =5$\ast$=5∗ = 5 in (17) as the positive semidefinite matrix
    $\overline{G_{f}}=\begin{pmatrix}5&5&5&5\\ 5&5&5&5\\ 5&5&5&5\\ 5&5&5&5\\ \end{pmatrix}$
    is clearly a positive semidefinite completion of $G_{f}$. But, this doesn’t contradict Theorem 3.35.
    To see this, we first note that determinant of the matrix
    $\begin{pmatrix}5&5&\ast\\ 5&5&5\\ \ast&5&5\end{pmatrix}$ (18)
    is $-(\ast-5)^{2}$. Therefore, the matrix as in (18) is positive semidefinite if and only if =5$\ast$=5∗ = 5. That is to say, the only positive semidefinite completion of this partial positive semidefinite matrix is
    $\begin{pmatrix}5&5&5\\ 5&5&5\\ 5&5&5\end{pmatrix}\;.$
    We now exploit this uniqueness. Consider the polynomial
    $\begin{split}f^{\prime}=&5\zeta_{1}^{\ast}\zeta_{1}+5\zeta_{2}^{\ast}\zeta_{2}+5\zeta_{3}^{\ast}\zeta_{3}+5\zeta_{4}^{\ast}\zeta_{4}+5(\zeta_{1}^{\ast}\zeta_{2}+\zeta_{2}^{\ast}\zeta_{1})+\ast(\zeta_{1}^{\ast}\zeta_{3}+\zeta_{3}^{\ast}\zeta_{1})\\ &+\sqrt{5}(\zeta_{1}^{\ast}\zeta_{4}+\zeta_{4}^{\ast}\zeta_{1})+5(\zeta_{2}^{\ast}\zeta_{3}+\zeta_{3}^{\ast}\zeta_{2})+\ast(\zeta_{2}^{\ast}\zeta_{4}+\zeta_{4}^{\ast}\zeta_{2})+5(\zeta_{3}^{\ast}\zeta_{4}+\zeta_{4}^{\ast}\zeta_{3})\\ =5&+(5+\ast)X_{1}X_{2}+(5+\ast)X_{2}X_{1}+2\sqrt{5}X_{2}^{2}+5X_{1}X_{2}^{2}X_{1}+5X_{2}X_{1}^{2}X_{2}+5X_{1}X_{2}X_{1}X_{2}\\ &+5X_{2}X_{1}X_{2}X_{1}+5X_{1}X_{2}^{3}+5X_{2}^{3}X_{1}+\ast X_{2}X_{1}X_{2}^{2}+\ast X_{2}^{2}X_{1}X_{2}+5X_{2}^{4}.\end{split}$ (19)
    Again it is a qusi SOHS polynomial of degree $4$ and associated partial positive semidefinite Gram-like matrix $G_{f^{\prime}}$ is given by
    $G_{f^{\prime}}=\begin{pmatrix}5&5&\ast&\sqrt{5}\\ 5&5&5&\ast\\ \ast&5&5&5\\ \sqrt{5}&\ast&5&5\\ \end{pmatrix}$ (20)
    having specification graph $\mathcal{C}_{4}$. Suppose, the matrix $G_{f^{\prime}}$ is positive semidefinite completable. Then first consider the leading principal submatrix of $G_{f^{\prime}}$ obtained by deleting last row and column. Then, as this has to be positive semidefinite after completion, $\ast$ in $(1,3)$-th and $(3,1)$-th position must have to be $5$ by the aforementioned uniqueness. Now, considering the principal submatrix obtained by deleting first row and column and applying same argument, we obtain $\ast$ in $(2,4)$-th and $(4,2)$-th position must have to be $5$. But then the principal submatrix of $G_{f^{\prime}}$, obtained either by deleting second row and second column (or by deleting third row and third column), has determinant $(\sqrt{5}-5)(25-5\sqrt{5})(<0)$. Therefore, the matrix $G_{f^{\prime}}$ is not positive semidefinite completable and hence the polynomial, as in (19) doesn’t have a SOHS completion.
We now go through some applications of Theorem 3.35.
Proposition 3.37.
Let $f$ be a quasi SOHS polynomial of degree $2d$ w.r.t to the representation $\mathbf{R}_{f}$
$\mathbf{R}_{f}\;:\;f=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$
as in (8) along with conditions as in (9). If out of $\tfrac{k(k-1)}{2}$ many $b_{ij}$s, any $k-1$ are specified and remaining $\tfrac{(k-1)(k-2)}{2}$ are unspecified, and if there is no column (or row) of the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$ with diagonal entry as the only specified entry, then $f$ has a SOHS completion.
Proof.
By hypothesis, in any column of $G_{f}$ other than the diagonal entry at least one entry is specified. This ensures that the specification graph $\mathcal{G}_{G_{f}}$ of $G_{f}$ is connected. Moreover, the hypothesis that any $k-1$ many $b_{ij}$s are specified affirms that there are exactly $k-1$ edges between $k$ many vertices of $\mathcal{G}_{G_{f}}$. Therefore, the graph $\mathcal{G}_{G_{f}}$ is indeed a tree and hence chordal. Hence, the assertion follows from Theorem 3.35. ∎
Corollary 3.38.
Let $f$ be a quasi SOHS polynomial of degree $2d$ w.r.t to the representation $\mathbf{R}_{f}$
$\mathbf{R}_{f}\;:\;f=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$
as in (8) along with conditions as in (9). If out of $\tfrac{k(k-1)}{2}$ many $b_{ij}$s, exactly $k-1$ are specified for some fixed $i$ (or fixed $j$), then $f$ has a SOHS completion.
Proof.
Follows from Proposition 3.37. ∎
Remark 3.39.
It can be noted that, taking $k=4$ in Corollary 3.38, the specification graphs $\mathcal{G}_{G_{f}}$ are the following rooted trees of height one on $4$ vertices.
[…]
Corollary 3.40.
Let $f$ be a quasi SOHS polynomial of degree $2d$ w.r.t to the representation $\mathbf{R}_{f}$
$\mathbf{R}_{f}\;:\;f=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$
as in (8) along with conditions as in (9). If out of $\tfrac{k(k-1)}{2}$ many $b_{ij}$s, at most $3$ are specified, then $f$ has a SOHS completion.
Proof.
The hypothesis ensures that the specification graph $\mathcal{G}_{G_{f}}$ of the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$ has $k$ vertices and at most $3$ edges. Therefore, $\mathcal{G}_{G_{f}}$ can’t have $\mathcal{C}_{n}$ as an induced subgraph for $n\geq 4$ and $\mathcal{G}_{G_{f}}$ is chordal. Therefore it follows from Theorem 3.35. ∎
3.4. Positive semidefinite completion problem and SOHS via algebraic sets
In Subsection 3.3, we answered Question 1 using techniques of solving positive semidefinite completion problem via graph theory. Recall that, given a partial positive semidefinite matrix $G$, we considered a graph associated with it which was obtained by keeping track of the specified off diagonal entries of $G$. On the contrary, now we answer Question 1 using another method of solving positive semidefinite completion problem. We do so by associating a projective algebraic set to it that keeps track of the unspecified off diagonal entries. In the process, we obtain a nice connection between noncommutative polynomial and usual (commutative) polynomial.
In this section, by a polynomial we mean the usual commutative polynomial. We denote commutative variables by small alphabets, unlike the noncommutative variables. By a form we mean a homogeneous polynomial. It is obvious that if a (homogeneous) polynomial has a sum of square (SOS) representation then it should be nonnegative. But the converse is not true. In fact, Hilbert, in 1888, proved that converse is true only for quadratic forms, binary forms and ternary quartics (cf. [19]). Many years later, Motzkin gave a specific example of ternary sextic, namely $x_{0}^{6}+x_{1}^{4}x_{2}^{2}+x_{1}^{2}x_{2}^{4}-3x_{0}^{2}x_{1}^{2}x_{2}^{2}$, which is a nonnegative polynomial but can not be expressed as a SOS. In 2016, Blekherman, Smith and Velasco has generalised Hilbert’s result. Instead of looking into globally nonnegative polynomials, they looked at nonnegative quadratic forms over a nondegenerate totally-real projective variety $X\subset\mathbb{P}^{n}$. Here, by a totally-real variety they meant that the set of real points in $X$ is a dense subset of set of complex points of $X$. They proved that any nonnegative polynomial would be a sum of squares, modulo the ideal of $X$ if and only if $X$ is a variety of minimal degree, i.e, ${\rm deg}(X)={\rm codim}(X)+1$ (cf. [8, Theorem 1.1,p. 893]). In 2017, Blekherman, Sinn and Velasco generalised it even further for any nondegenerate totally-real projective algebraic set. As the notion of degree doesn’t make sense for reducible algebraic sets, they observed that a right generalisation of the notion of degree is Castelnuovo-Mumford regularity. We now recall the meaning of this notion for a projective algebraic set.
Definition 3.41.
Let $k$ be a field of characteristic 00. A finitely generated $k[x_{0},x_{1},\ldots,x_{n}]$-module $M$ has a unique free resolution
$\cdots\rightarrow F_{t}\rightarrow F_{t-1}\rightarrow\cdots\rightarrow F_{0}\rightarrow M\rightarrow 0.$
Let $b_{j}$ be the largest degree of a minimal generator in the module $F_{j}$. We also fix the convention that $b_{j}=-\infty$ if $F_{j}$ is the zero module. Then the module $M$ is said to be $m$-regular if $b_{j}\leq m+j$ for all $j$.
Definition 3.42.
A projective algebraic set $X\subset\mathbb{P}^{n}$ is said to be $m$-regular if the defining ideal $I$ of $X$ is $m$-regular as a $k[x_{0},x_{1},\ldots,x_{n}]$-module.
We now state the main result obtained by Blekherman, Sinn and Velasco generalising [8, Theorem 1.1,p. 893].
Theorem 3.43.
Let $X\subset\mathbb{P}^{n}$ be a totally-real projective algebraic set. Every nonnegative quadratic form on $X$ is a sum of squares, modulo the ideal of $X$ if and only if $X$ is 2-regular.
Proof.
See [6, Theorem 9, p. 181]. ∎
Remark 3.44.
It can be noted that linear projective varieties are $1$-regular. Apart from that, any varieties of minimal degree are $2$-regular. That is why, the apt generalisation of the notion of degree for a reduced algebraic set is regularity.
Apparently, it might seem that the results proved in [8] and [6] are only for quadratic forms, whereas Hilbert proved the same for any $2d$ degree polynomials. But, they showed that it is enough to consider quadratic forms as any other even degree forms can be settled down using Veronese embedding. To be precise, to look at degree $2d$ forms on a variety $X$, it is enough to consider quadratic forms on the image $\nu_{d}(X)$ of $X$ via $d$-th Veronese embedding $\nu_{d}$, (cf. [7, p. 737]).
Both of the results, i.e, Theorem 1.1 in [8] and Theorem 9 in [6], have a deep connection with positive semidefinite completion problem. Recall that, given a partial positive semidefinite matrix $G$, its specification graph encodes the known part of $G$. On the contrary, to encode the unknown part of $G$ a projective algebraic set can be associated with it. In this regard, we have the following definition.
Definition 3.45.
The subspace arrangement $V_{A}$ of a $n\times n$ partial symmetric matrix $A$ is a projective algebraic subset of $\mathbb{P}^{n-1}$ whose vanishing homogeneous ideal is generated by the (square free) quadratic monomials $x_{i-1}x_{j-1}$, $1\leq i<j\leq n$, whenever $(i,j)$-th entry of $A$ is unspecified.
As mentioned earlier, it is expected that given a partial symmetric matrix, its specification graph and subspace arrangement should be related. Following result by Fröberg points out that relation. In the process, it provides a bridge between algebraic geometry, commutative algebra, graph theory and positive semidefinite completion problem.
Theorem 3.46.
Let $G$ be a partial symmetric matrix. Then the subspace arrangement $V_{G}$ is $2$-regular if and only if the specification graph $\mathcal{G}_{G}$ of $G$ is chordal.
Proof.
See [13]. ∎
Unlike commutative setup, the relation between positive noncommutative polynomials and SOHS polynomials is simple. In fact, any positive noncommutative polynomial is a SOHS polynomial (cf. Remark 2.3). But the equality between the nonnegative quadratic forms and SOS modulo a suitable projective algebraic set has the potential to allow a SOHS extension of a quasi SOHS polynomial of any even degree. We encode this in the following theorem, which is nothing but a reinterpretation of Theorem 3.35 through the lens of algebraic geometry and commutative algebra. Before that, we quickly recall a couple of definitions along with their notations.
We denote by $S$ the homogeneous coordinate ring of a projective algebraic set $X\subset\mathbb{P}^{n}$. That is, if $I$ is the homogeneous ideal of $\mathbb{R}[x_{0},\ldots,x_{n}]$ generated by all forms that vanish on $X$, then $S:=\tfrac{\mathbb{R}[x_{0},\ldots,x_{n}]}{I}$. By $S_{j}$ we denote the vector space over $\mathbb{R}$ spanned by the degree $j$ forms of $S$. Then by $P_{X}$ and $\Sigma_{X}$ we denote the following two sets.
[…]
Theorem 3.47.
Let $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ be a quasi SOHS polynomial w.r.t a representation $\mathbf{R}_{f}$, as in (8), along with conditions as in (9). Then the following are equivalent:
  1. (1)
    Any such $f$ has a SOHS completion $\overline{f}^{\;G_{f},\;W_{k}}$ w.r.t the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$.
  2. (2)
    The subspace arrangement $V_{G_{f}}(\subseteq\mathbb{P}^{k-1})$ of $G_{f}$ satisfies $P_{V_{G_{f}}}=\Sigma_{V_{G_{f}}}$.
  3. (3)
    The subspace arrangement $V_{G_{f}}(\subseteq\mathbb{P}^{k-1})$ of $G_{f}$ is $2$-regular.
Here $W_{k}$ is the column vector $\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}\end{pmatrix}^{t}$, $\zeta_{i}$, $1\leq i\leq k$, being the monomials appearing in $\mathbf{R}_{f}$.
Proof.
Follows from Theorem 3.35, Theorem 3.43 and Theorem 3.46. ∎
We now apply Theorem 3.35 and Theorem 3.47 to provide a couple of examples of some quasi SOHS polynomials that admit SOHS completions.
Example 3.48.
Consider any quasi SOHS polynomial $f$ whose partial positive semidefinite Gram-like matrix $G_{f}$ is given as follows :
$G_{f}=\begin{pmatrix}5&5&5&\ast\\ 5&5&5&\ast\\ 5&5&5&5\\ \ast&\ast&5&5\\ \end{pmatrix}.$
Then the corresponding subspace arrangement $V_{G_{f}}$ is the projective algebraic set sitting inside $\mathbb{P}^{3}$ whose defining ideal is the homogeneous ideal $I:=\langle x_{0}x_{3},x_{1}x_{3}\rangle$. Using Macaulay2 software (cf. [14]), we compute the free resolution and Betti table of $I$, considered as a $\mathbb{R}[x_{0},x_{1},x_{2},x_{3}]$-module, as follows :
Free resolution of $I$ : Denoting $\mathbb{R}[x_{0},x_{1},x_{2},x_{3}]$ by $R$, we have :
[…]
Betti table of $I$:
0   1total :2   12:2   $\begin{split}&0\;\;\;1\\ \text{total :}\;&2\;\;\;1\\ 2:\;&2\;\;\;1\end{split}$
Therefore, we obtain that the ideal $I$ is $2$-regular and so is the projective algebraic set $V_{G_{f}}$. Therefore, by Theorem 3.47, any such quasi SOHS polynomial $f$ admits a SOHS completion w.r.t $G_{f}$. This is also evident from Theorem 3.35 as the specification graph $\mathcal{G}_{G_{f}}$
[…]
is chordal.
Example 3.49.
Consider any quasi SOHS polynomial $f$ whose partial positive semidefinite Gram-like matrix $G_{f}$ has the following disconnected specification graph $\mathcal{G}_{f}$ :
[…]
Then the corresponding subspace arrangement $V_{G_{f}}$ is the projective algebraic subset of $\mathbb{P}^{4}$ whose defining ideal is the homogeneous ideal $I:=\langle x_{0}x_{2},x_{0}x_{3},x_{0}x_{4},x_{1}x_{2},x_{1}x_{3},x_{1}x_{4},x_{2}x_{4},x_{3}x_{4}\rangle$. Using Macaulay2 software, we compute the free resolution and Betti table of $I$ considered as a $\mathbb{R}[x_{0},x_{1},x_{2},x_{3},x_{4}]$-module, as follows :
Free resolution of $I$ : Denoting $\mathbb{R}[x_{0},x_{1},x_{2},x_{3},x_{4}]$ by $R$, we have :
[…]
Betti table of $I$:
0    1    2   3total :8   14   9   22:8   14   9   $\begin{split}&0\;\;\;\;1\;\;\;\;2\;\;\;3\\ \text{total :}\;&8\;\;\;14\;\;\;9\;\;\;2\\ 2:\;&8\;\;\;14\;\;\;9\;\;\;2\\ \end{split}$
Therefore, we obtain that the ideal $I$ is $2$-regular and so is the projective algebraic set $V_{G_{f}}$. Therefore, by Theorem 3.47, any such quasi SOHS polynomial admits a completion w.r.t $G_{f}$. In fact, it also follows from Theorem 3.35 as the specification graph $\mathcal{G}_{f}$ is chordal.
We now apply Theorem 3.35 and Theorem 3.47 to provide an example of some quasi SOHS polynomials that do not admit SOHS completions.
Example 3.50.
Consider the partial positive semidefinite matrix, as in (20),
$G^{\prime}=\begin{pmatrix}5&5&\ast&\sqrt{5}\\ 5&5&5&\ast\\ \ast&5&5&5\\ \sqrt{5}&\ast&5&5\\ \end{pmatrix}.$
Then the corresponding subspace arrangement $V_{G^{\prime}}$ is the projective algebraic set sitting inside $\mathbb{P}^{3}$ whose defining ideal is the homogeneous ideal $J:=\langle x_{0}x_{2},x_{1}x_{3}\rangle$. Using Macaulay2 software (cf. [14]), we compute the free resolution and Betti table of $I$, considered as a $\mathbb{R}[x_{0},x_{1},x_{2},x_{3}]$-module, as follows :
Free resolution of $J$ : Denoting $\mathbb{R}[x_{0},x_{1},x_{2},x_{3}]$ by $R$, we have :
[…]
Betti table of $J$:
0   1total :2   12:2.3:.   $\begin{split}&0\;\;\;1\\ \text{total :}\;&2\;\;\;1\\ 2:\;&2\;\;\;\textbf{.}\\ 3:\;&\textbf{.}\;\;\;1\\ \end{split}$
Therefore, we obtain that the ideal $J$ is $3$-regular and so is the projective algebraic set $V_{G^{\prime}}$. Therefore, by Theorem 3.47, there exists a quasi SOHS polynomial that doesn’t admit a SOHS completion w.r.t $G^{\prime}$. In fact, the polynomial $f^{\prime}$, as given in (19), is one such. This is not all surprising. In fact, it follows from Theorem 3.35 as the specification graph $G^{\prime}$
[…]
is the cycle $\mathcal{C}_{4}$ on four vertices, which is not a chordal graph.
Remark 3.51.
The flexibility of a Gram-like matrix over a Gram matrix is quite evident from these examples. As we deal with Gram-like matrices, we could play around with graphs having any number of vertices. For a Gram matrix, we can only deal with graphs with $s(d,n)$ many vertices for some $d\geq 0$ and $n>0$.
Remark 3.52.
As a next step, it would be interesting to generate algorithms out of the criteria provided in Theorem 3.19, Theorem 3.22, Theorem 3.24, Theorem 3.35 and Theorem 3.47 and use that further to produce SDP corresponding to these extension problems.
Acknowledgements
Both the authors extend their gratitude to Prof. Sumesh K, Prof. Arijit Dey and Prof. Sriram B for their valuable suggestions regarding the article. The authors acknowledge Rajib Sarkar and Sreenanda S B for their assistance regarding the use of Macaulay2 software. The authors thank Dr. Aditi Howlader for suggesting some useful references. The authors also thank Agnihotri Roy for reexamining some tedious calculations. The first named author would like to thank Indian Institute of Technology Madras for financial support (Office order No.F.ARU/R10/IPDF/2024). The second named author would like to thank Indian Institute of Technology Madras and ICSR for financial support (Project number - SB22231267MAETWO008573).